### 04 Apr 2026 v1

# Magnitude homology of real hyperplane arrangements

Abstract.
We initiate the study of magnitude theory of real hyperplane arrangements. Magnitude is a cardinality-like invariant of metric spaces or enriched categories measuring the effective size. Its categorification, the magnitude homology, is a more powerful invariant. For a real hyperplane arrangement, or more generally, an oriented matroid, the tope graph encapsulates considerable amount of information. Since tope graphs are equipped with the shortest path metric, we feed them to the magnitude and magnitude homology machinery to derive new invariants of real hyperplane arrangements. We prove some structural results of the magnitude of arrangements, including reciprocity, palindromic numerator and denominator. For magnitude homology of arrangements, we give combinatorial descriptions in small length and prove that tope graphs are diagonal if and only if the arrangement is Boolean. We present a face decomposition of magnitude homology, using which we obtain a combinatorial formula of diagonal magnitude Betti numbers. Many open problems are posted for future study. In particular, we conjecture that magnitude and magnitude homology of arrangements are determined by the intersection lattice.
Key words and phrases:
hyperplane arrangement, tope graph, magnitude, magnitude homology
2020 Mathematics Subject Classification:
Primary 52C35; Secondary 55N35
1. Introduction
Let $\mathcal{A}$ be an arrangement of hyperplanes in $\mathbb{R}^{\ell}$ all passing through the origin. The hyperplanes in $\mathcal{A}$ cut $\mathbb{R}^{\ell}$ into a disjoint union of chambers. Let $\mathrm{Ch}(\mathcal{A})$ denote the set of chambers. For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ as the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$.
The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$, and two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane. The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$. Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, isometric to a subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ equipped with the Hamming distance. For an oriented matroid $\mathcal{M}$, the tope graph $\mathcal{T}(\mathcal{M})$ can be defined similarly.
Although seemingly elementary, the tope graphs turn out to be quite informative.
Theorem 1.1 (Theorem 4.2.14 of [4]).
A simple oriented matroid is uniquely determined (upto reorientation) by its unlabeled tope graph.
We refer to Figure 1 of [32] for a detailed diagram of implications. In particular, if two arrangements have isomorphic tope graphs, then they have homoemorphic complexified complements, isomorphic intersection lattices and isomorphic filtered Varchenko–Gelfand algebras. Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
In this paper, we feed tope graphs into the emerging machinery of magnitude theory. Originally introduced by Leinster for enriched categories and metric spaces [24, 21], magnitude acts as a cardinality-like invariant measuring the “effective size” of a space. For finite graphs [25], it is a rational function defined via the inverse of the $q$-distance matrix, which is intimately related to the classical Varchenko matrix in the context of arrangements. The categorification of this concept, magnitude homology, developed by Hepworth and Willerton for graphs [16], and by Leinster and Shulman for enriched categories and metric spaces [23], provides a powerful bigraded homology theory that detects profound metric features.
We initiate the study of magnitude and magnitude homology specifically tailored to the metric geometry of real hyperplane arrangements. Our motivation is two-fold. The first is that we introduce a new rational function invariant and a new bigraded homology invariant of real arrangements. As a special class of graphs, tope graphs of arrangements (or oriented matroids) possess many unique features, such as antipodal symmetry and organized face decomposition, that enable us to contribute to the theory of magnitude with new beautiful examples.
At the decategorified level, we demonstrate that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, exhibits remarkable structural results. We prove a reciprocity formula relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, and establish that its reduced rational form consists of palindromic numerator and denominator whose poles are roots of unity. Supported by extensive computational evidence generated via SageMath on the Grünbaum catalogue [14], we further propose a conjectural Face Decomposition formula, which serves as a discrete $q$-analogue to the classical decomposition of the Euler characteristic of a zonotope.
At the homological level, by exploiting the metric structure of gated subgraphs within partial cubes, which naturally arise from the faces of the arrangement, we systematically dissect the magnitude homology of $\mathcal{T}(\mathcal{A})$. After providing explicit combinatorial descriptions of the homology groups for small lengths, we establish a strict topological dichotomy for arrangements. We prove that the tope graph is diagonal, meaning its off-diagonal magnitude homology strictly vanishes, if and only if the arrangement is Boolean. We also prove a face decomposition of magnitude homology of arrangements, and as an application, we derive a formula of diagonal magnitude Betti numbers in terms of the intersection lattice.
The paper is organized as follows. Section 2 collects necessary recollections from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition. Section 3 investigates the basic properties and structural rigidity of the magnitude rational function. Section 4 explores magnitude homology, presenting the Boolean diagonality theorem and face decomposition theorem. Finally, Section 5 posts several open problems and conjectures for future study. Results of many computer experiments with SageMath form the appendix.
Convention: When we talk about a topological property/construction of a poset, we mean that of its order complex.
2. Recollections
2.1. Metric graph theory
In this paper, all graphs we mention are supposed to be finite simple undirected graphs unless otherwise specified. For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from vertex $x\in V$ to vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$ and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$. We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ is the minimal length of walks from $x$ to $y$. If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$. Then $(G,d)$ is a generalized metric space. In particular, when $G$ is connected, that is there is a walk between any pair of vertices, $(G,d)$ is a metric space. We call the maximal value of $d$ the diameter of $G$, denoted by $\mathrm{diam}(G)$.
A subgraph $H$ of $G$ can be equipped with a metric in two ways. One is to restrict $d_{G}$ to $H$ and the other is $d_{H}$ on $H$ as a graph itself. We say $H$ is an isometric subgraph of $G$ if the two metrics coincide $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk. Given $x,y\in V$, let us define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$. It is a poset equipped with partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ traversing $u$ before traversing $v$. Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$. A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$. More generally, vertices $u_{1},\ldots,u_{k}$ form a strict linearly ordered chain in $x,y]_{G}$, that is $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$ if and only if
$d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ excluding $x$ and $y$.
We say a subgraph $H$ of $G$ is convex in $G$ if for any $x,y\in V(H)$, any minimal walk from $x$ to $y$ traverses only vertices in $H$, equivalently $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$. Apparently, any convex subgraph is an isometric subgraph, but the converse is false.
Definition 2.1.
Let $G$ be a connected graph, $H$ an isometric subgraph of $G$ and $x\in V(G)$. We say $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that for any $v\in V(H)$,
$d(x,v)=d(x,u)+d(u,v).$
Such $u$ must be unique since the above equation says $u$ is the closest vertex to $x$ in $H$. We may call $u$ the gate of $H$ with respect to $x$. Intuitively, any minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate. We say $H$ is a gated subgraph of $G$ is it is gated with respect to all vertices of $G$. In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection. Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
Proposition 2.2 (Proposition A.1 of [2]).
A gated subgraph is necessarily convex.
Proof.
Suppose $H$ is gated in $G$. Choose any minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$ and let $z$ be an arbitrary vertex in this walk, we need to show $z\in V(H)$. Since $H$ is gated, we have
d(z,x)=d(z,p(z))+d(p(z),x),d(z,y)=d(z,p(z))+d(p(z),y).d(z,x)=d(z,p(z))+d(p(z),x),~d(z,y)=d(z,p(z))+d(p(z),y).
Since $z$ belongs to $x,y]_{G}$, adding these two equations yields
$d(x,y)=d(x,z)+d(z,y)=d(x,p(z))+d(p(z),y)+2d(z,p(z))\geq d(x,y)+2d(z,p(z)).$
This forces $d(z,p(z))=0$, i.e. $z=p(z)\in V(H)$. ∎
In the scope of this paper, the homomorphisms between graphs we shall use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, which are maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$. Note that gate projections are distance non-increasing.
2.2. Graph magnitude
See Leinster [25] for details. For a finite simple graph $G=(V,E)$, define the Zeta matrix $Z_{G}=Z_{G}(q)=[q^{d(x,y)}]_{x,y\in V}$, where $q^{\infty}=0$ by convention. Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$. The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ then can be expressed as a power series in $\mathbb{Z}[\![q]\!]$. In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
The magnitude of $G$ is defined by the sum of all entries of $Z_{G}(q)^{-1}$.
$\mathrm{Mag}(G)=\mathrm{Mag}(G,q):=\sum_{x,y\in V}Z_{G}(q)^{-1}_{x,y}\in\mathbb{Q}(q)~(\mathrm{or}~\mathbb{Z}[\![q]\!]).$
In general, the sum of all entries of a matrix $A$ is equal to $\mathbf{1}^{T}A\mathbf{1}$, where $\mathbf{1}$ is the column vector of appropriate size with all entries equal to $1$. If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$. For graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$, in other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$. Then the magnitude
$\mathrm{Mag}(G)=\mathbf{1}^{T}w_{G}=\sum_{x\in V}w_{G}(x).$
There is a simple case that we can compute magnitude easily.
Proposition 2.3 (Lemma 3.2 of [25]).
If $G$ is vertex-transitive (i.e. $\mathrm{Aut}(G)$ acts transitively on $V$), then
$\mathrm{Mag}(G)=\frac{\#V}{\sum_{x\in V}q^{d(g,x)}},$
for any $g\in V$.
Proof.
The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and thus denoted by $s=s(g)$. Then it is easily seen that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$. We conclude $\mathrm{Mag}(G)=\#V/s$. ∎
Example 2.4.
Let $K_{n}$ be the complete graph on $n$ vertices and $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Then
$\mathrm{Mag}(K_{n})$ $=\frac{n}{1+(n-1)q},$
$\mathrm{Mag}(C_{2n})$ $=\frac{2n(q-1)}{(q^{n}-1)(q+1)}=\frac{2n}{(1+q)(1+q+q^{2}+\cdots+q^{n-1})},$
$\mathrm{Mag}(C_{2n-1})$ $=\frac{(2n-1)(q-1)}{2q^{n}-q-1}=\frac{2n-1}{1+2q+2q^{2}+\cdots+2q^{n-1}}.$
Magnitude behaves as cardinality in the following sense.
Proposition 2.5 (Lemmas 3.5 and 3.6 of [25]).
Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ the disjoint union and cartesian product of $G$ and $H$ respectively. Then
$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),~\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
Under some conditions, magnitude satisfies inclusion-exclusion.
Theorem 2.6 (Theorem 4.9 of [25]).
If a graph $X$ has a decomposition into subgraphs $X=G\cup H$ such that $G\cap H$ is convex in $X$ and gated in $H$, then
$\mathrm{Mag}(X)=\mathrm{Mag}(G)+\mathrm{Mag}(H)-\mathrm{Mag}(G\cap H).$
2.3. Magnitude homology of graphs
For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$. We say a $k$-chain $\vec{x}$ is proper if adjacent nodes are different $x_{0}\neq x_{1}\neq\cdots\neq x_{k}$. Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and $P_{k,\ell}(G)$ the set of proper $k$-chains of length $\ell$. The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [16].
Proposition 2.7 (Proposition 3.9 of [25]).
For a graph $G$,
$\mathrm{Mag}(G)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{\vec{x}\in P_{k}(G)}q^{\ell(\vec{x})}\in\mathbb{Z}[\![q]\!].$
In other words, if $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$, then
$c_{\ell}=\sum_{k=0}^{\ell}(-1)^{k}\#P_{k,\ell}(G).$
In particular, $V=c_{0}$ and $E=\frac{-c_{1}}{2}$.
Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$. We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ as $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
$\partial_{i}(x_{0},\ldots,x_{k})=\begin{cases}(x_{0},\ldots,\widehat{x_{i}},\ldots,x_{k}),&\text{ if }d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1});\\ 0,&\text{ otherwise},\end{cases}$
where $\widehat{x_{i}}$ means deletion of $x_{i}$. One can verify $\partial^{2}=0$. Therefore we call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$. Since the boundary map preserves endpoints, we have a decomposition
$MC_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MC_{k,\ell}(G;u,v),$
where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,x_{k}=v\}$, forming subcomplexes and inducing decomposition of homology groups
(2.1) $MH_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MH_{k,\ell}(G;u,v).$
Let us denote the magnitude Betti number $\beta_{k,\ell}(G)=\mathrm{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
Proposition 2.8 (Propositions 2.9, 2.10 of [16]).
Let $G=(V,E)$ be a graph.
  1. (1)
    $MH_{0,0}(G)$ is free abelian of rank $\beta_{0,0}(G)=\#V$.
  2. (2)
    $MH_{1,1}(G)$ is free abelian of rank $\beta_{1,1}(G)=2\#E$.
  3. (3)
    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\mathrm{diam}(G)\leq k\leq\ell$ and the inequality $\ell/\mathrm{diam}(G)<k$ is strict if $\mathrm{diam}(G)>1$ and $\ell>0$.
The following proposition is straightforward from definition, which says magnitude can be recovered from magnitude homology.
Proposition 2.9 (Theorem 2.8 of [16]).
Let $G$ be a graph, and $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$. Then
$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\mathrm{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
Hepworth–Willerton proved that magnitude homology is a functor.
Proposition 2.10 (Proposition 3.3 of [16]).
Magnitude homology is a functor from the category of finite graphs with distance non-increasing maps to the category of bigraded abelian groups.
A direct application of the functorality is as follow. Let $G$ be a connected graph and $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$. Let $\iota:H\to G$ be the inclusion. Both $p$ and $\iota$ are distance non-increasing. The composition $p\circ\iota=\text{id}_{H}$ induces $\text{id}=p_{*}\circ\iota_{*}:MH_{k,\ell}(H)\to MH_{k,\ell}(G)\to MH_{k,\ell}(H)$. This gives a decomposition of $MH_{k,\ell}(G)$.
Proposition 2.11.
If $p:G\to H$ is a gate projection, then
$MH_{k,\ell}(G)\cong MH_{k,\ell}(H)\oplus\ker(p_{*}).$
There is a Künneth formula for magnitude homology of graphs.
Theorem 2.12 (Theorem 5.3 of [16]).
There is a natural short exact sequence which non-naturally splits
$0\to MH_{*,*}(G)\otimes MH_{*,*}(H)\xrightarrow{\square}MH_{*,*}(G\square H)\to\mathrm{Tor}(MH_{*-1,*}(G),MH_{*,*}(H))\to 0,$
where $\square$ is the exterior product defined in Definition 5.2 of [16].
Example 2.13.
For the complete graph $K_{n}$,
$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=l;\\ 0,&\text{ otherwise}.\end{cases}$
The Betti number $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
Example 2.14.
For the hypercube graph $Q_{d}=K_{2}^{\square d}$, we use the Künneth formula (Theorem 2.12) to obtain
$MH_{k,\ell}(Q_{d})=\bigoplus_{\sum k_{i}=k,\sum\ell_{i}=\ell}MH_{k_{1},\ell_{1}}(K_{2})\otimes\cdots\otimes MH_{k_{d},\ell_{d}}(K_{2}),$
and use the result of $K_{2}$, we conclude
$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=l;\\ 0,&\text{ otherwise},\end{cases}$
where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{l+d-1}{d-1}$. One can verify $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
Definition 2.15.
We say $G$ is diagonal if $MH_{k,\ell}(G)=0$ for $k\neq\ell$.
Example 2.16.
We have seen that $K_{n}$ and $Q_{d}$ are diagonal. In particular the cycle graph $C_{4}=Q_{2}$ is diagonal. However Gu [15] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that $C_{n}(n\geq 5)$ are not diagonal.
Let us recall results from [19] in the context of graphs. Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$. We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$. Any sub-chain of a geodesic proper chain is again geodesic proper. Any proper $1$-chain is geodesic. A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$.
Consider $u,v\in V(G)$, such that $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
Definition 2.17.
Let us define the length $\ell$ geodesic magnitude complex
$MC_{*,\ell}^{geod}(G):=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MC_{*,\ell}(G;u,v).$
The homology is called the length $\ell$ geodesic magnitude homology, which admits the decomposition
$MH_{k,\ell}^{geod}(G)=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MH_{k,\ell}(G;u,v).$
Note that $MC_{*,\ell}^{geod}(G)$ and $MH_{k,\ell}^{geod}(G)$ are trivial if $\ell>\mathrm{diam}(G)$.
Let us denote the geodesic magnitude Betti number $\beta_{k,\ell}^{geo}(G)=\mathrm{rank}MH_{k,\ell}^{geod}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic $\chi_{\ell}^{geod}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{geod}(G)$. Notice that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{geod}(G)$.
Kaneta–Yoshinaga [19] generalized the above geodesic magnitude chains.
Definition 2.18.
A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic. Write $m_{G}$ as the minimum length of $4$-cuts of $G$. We say $m_{G}=\infty$ if $G$ has no $4$-cuts.
For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}(i=1,2,\ldots,k-1)$ is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$. A node that is not smooth is called a singular node. Note that endpoints are singular. We write $\varphi(\vec{x})$ as the sub-chain of $\vec{x}$ consisting of all singular nodes, called the frame of $\vec{x}$. Note that $\varphi(\vec{x})$ is not necessarily proper.
Definition 2.19.
A proper chain $\vec{x}$ is geodesically simple if $\ell(\varphi(\vec{x}))=\ell(\vec{x})$.
Note that proper $1$-chains and proper $2$-chains are all geodesically simple, and proper chains with length smaller than $m_{G}$ are geodesically simple. Kaneta–Yoshinaga proved that the subgroups $MC_{*,\ell}^{gs}(G)$ of $MC_{*,\ell}(G)$ generated by geodesically simple chains is a subcomplex, whose homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{gs}(G)$, and the homology group $MH_{k,\ell}^{gs}(G)$ factors as a direct sum of framed magnitude homology groups as explained below.
Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$. We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$. Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and would be empty if $\varphi(\vec{a})\neq\vec{a}$. We further define $MC_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{gs}(G)$. They form a subcomplex $MC_{*,\vec{a}}(G)$ whose homology $MH_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$. Kaneta–Yoshinaga proved the chain complex decomposes
(2.2) $MC_{k,\ell}^{gs}(G)=\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MC_{k,\vec{a}}(G),$
and so does the homology group.
Theorem 2.20 (Theorem 3.12 of [19]).
Let $G$ be a graph and $k,\ell>0$. We have
$MH_{k,\ell}^{gs}(G)\cong\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MH_{k,\vec{a}}(G).$
Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{gs}(G)\cong MH_{k,\ell}(G)$.
Kaneta–Yoshinaga also related framed magnitude homology with open intervals determined by the frame. Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1. Also write $C_{*}(a,b)_{G}$ the reduced chain complex of the order complex of $(a,b)_{G}$.
Proposition 2.21 (Proposition 2.3 of [13]).
For $a,b\in G$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
$MC_{*,\ell}(G;a,b)\to C_{*-2}(a,b)_{G},$
inducing an isomorphism of homology groups
$MH_{k,\ell}(G;a,b)\xrightarrow{\cong}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
Proof.
For a proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$, it must be geodesic since $\ell=d(a,b)$. Then mapping $\vec{x}$ to the linearly order chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism. ∎
Corollary 2.22.
The geodesic magnitude homology has the decomposition
$MH_{k,\ell}^{geod}(G)\cong\bigoplus_{a,b\in V(G):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$. A more generalized version of Proposition 2.21 is as follow.
Theorem 2.23 (Theorem 4.4 and Corollary 4.5 of [19]).
Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$, then there is an isomorphism of chain complexes
$MC_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G}$
grouping smooth nodes in order between $a_{i}$ and $a_{i+1}$, which induces isomorphisms of homology groups
$MH_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
2.4. Hyperplane arrangements
In this paper, all arrangements are real central hyperplane arrangements unless otherwise specified. Standard references are [27, 30]. Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{\ell}$. We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusions. An element of $L(\mathcal{A})$ is called a flat. Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{\ell}$. The arrangement is called essential if the center is the origin. The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension $\mathrm{rank}X=\mathrm{codim}X$. The rank of $\mathcal{A}$ is defined as $\mathrm{rank}\mathcal{A}=\mathrm{rank}(\cap\mathcal{A})$. We write $L_{k}(\mathcal{A})$ as the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ the set of flats of rank at most $k$. Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$. It is classically known $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},1)|$ [33].
The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{\ell}$, the strata are called faces (covectors). Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusions. The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$. For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist. There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is the minimal subspace of $\mathbb{R}^{\ell}$ containing $F$. We define the dimension of a face as the dimension of its support.
For $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$. Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$. Note that if $F,G\in\mathcal{F}(\mathcal{A})$ have the same support $X\in L(\mathcal{A})$, then $\mathcal{A}_{F}=\mathcal{A}_{G}=\mathcal{A}_{X}$.
Definition 2.24.
For a flat $X\in L(A)$, we denote by
$c^{X}:=\#s^{-1}(X)$
the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}$$\mathbb{R}^{\ell}$, we also define the beta invariant of $X$ by
$\beta_{X}:=\beta(\mathcal{A}_{X}).$
Recall that the beta invariant of an arrangement is defined as
$\beta(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
Remark 2.25.
The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$. In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$, then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (See [33, 2, 30]). Hence the product $c^{X}\beta_{X}$ is also a counting function
$c^{X}\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid s(\overline{C}\cap H)=X\}.$
For arrangement $\mathcal{A}$ in $V=\mathbb{R}^{\ell}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$. We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ as the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$. The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we say the corresponding face $z_{F}$ of $Z(\mathcal{A})$ dual to $F$. In particular, 0-faces (vertices) of $Z(\mathcal{A})$ are dual to chambers of $\mathcal{A}$.
With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\subseteq\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$. Declare $0<\pm$ and this determines a partial order on sign vectors coordinate-wisely, which coincides with the face order of $\mathcal{F}(\mathcal{A})$. We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follow. For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(A)$ be the sign vector
$(FG)_{H}=\begin{cases}F_{H},\text{ if }F_{H}\neq 0;\\ G_{H},\text{ if }F_{H}=0.\end{cases}$
This product does not depend on the choices of defining forms, since geometrically, $FG$ is the face we land in when we start from a relative interior point of $F$, move out a small enough distance towards a relative interior point of $G$ (see Figure 1).
[…]
Figure 1. A geometric visualization of the non-commutative Tits product in the $A_{2}$ arrangement. Let $F$ be a ray (blue) and $G$ be a chamber (red). (a) The product FGFG represents taking a small step from $F$ towards $G$, which immediately lands in the adjacent chamber. (b) The product GFGF represents taking a small step from $G$ towards $F$. Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
2.5. Tope graphs and graph-theoretic properties
For an arrangement $\mathcal{A}$ in $\mathbb{R}^{\ell}$, notice that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$, or equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$. We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$. We first prove a useful lemma.
Lemma 2.26.
For $F\in\mathcal{F}(\mathcal{A})$, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$.
Proof.
We prove that the map $p_{F}:\mathcal{T}(\mathcal{A})\to\mathcal{T}(\mathcal{A})_{F}$ defined by $p_{F}(D)=FD$ is the gate projection, that is, we prove the equation
(2.3) $d(D,C)=d(D,FD)+d(FD,C)$
holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$. In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (number of non-equal signs), where
$d(X,Y)_{H}=\begin{cases}1,\text{ if }X_{H}\neq Y_{H};\\ 0,\text{ if }X_{H}=Y_{H}.\end{cases}$
Let us prove both sides of equation (2.3) have the same evaluation on each $H\in\mathcal{A}$. If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$, also since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$. Therefore (2.3) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is trivially true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$. Next suppose $H\in\mathcal{A}_{F}$, we have $F_{H}=0$. Then $d(D,FD)_{H}=0$. The equation (2.3) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is triviall true since $D_{H}=(FD)_{H}$. ∎
Here we record some important examples of tope graphs.
Example 2.27.
Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes of $\mathbb{R}^{d}$. It is clear that the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
Example 2.28.
Let $r\geq 2$. An arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$ is said to be generic if any subarrangement of $r$ hyperplanes has rank $r$. Boolean arrangements are generic. Generic arrangements are realizations of uniform matroids $U_{r,n}$. By abuse of notations, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$. For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$. The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ with all vertices (sign vectors) except the all ++ vector and the all - vector.
Example 2.29.
Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$. The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutohedron, and the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the Cayley graph of $(W,S)$.
Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if its intersection with the unit sphere of the ambient space is combinatorially a simplex. An arrangement is simplicial if all chambers all simplicial.
Proposition 2.30 (Propositions 4.2.15 and 4.4.8 of [4]).
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
  1. (1)
    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\mathrm{diam}(\mathcal{T})$.
  2. (2)
    The mapping $C\mapsto-C$ is a fixed point free automorphism of $\mathcal{T}$.
  3. (3)
    $\mathcal{T}$ is bipartite.
  4. (4)
    For each vertex $C$, $\deg(C)\geq r$, where the equality holds iff $C$ is simplicial.
  5. (5)
    $\mathcal{T}$ is $r$-connected.
Recall that a graph is called $k$-regular if every vertex has degree $k$.
Corollary 2.31.
$\mathcal{A}$ is simplicial of rank $r$ if and only if $\mathcal{T}(\mathcal{A})$ is $r$-regular.
A matching of a graph is a collection of edges where no two edges share a common incident vertex. We say a matching is perfect if every vertex of the graph is incident to exactly one edge in the matching. The marriage theorem (see Theorem 2.1.2 of [9]) has a corollary (see Corollary 2.1.3 of [9]) asserting that every $k$-regular ($k\geq 1$) bipartite graph has a perfect matching.
Corollary 2.32.
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A})$ has a perfect matching.
Very recently, Hamiltonian cycles in tope graphs of simplicial and supersolvable arrangements are studied in [5, 20].
2.6. Varchenko matrix
Varchenko introduced a bilinear form and computed the determinant [31]. Generalization to oriented matroids can be found in [17]. See [8] and Sections 8.4-8.5 of [2] for details.
Here we only consider the equal weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix
$\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]_{C,D\in\mathrm{Ch}(\mathcal{A})}.$
In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). It can also be defined for oriented matroids.
Theorem 2.33 ([31], see also Theorem 8.23 of [2]).
The determinant of the $q$-Varchenko matrix is
$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}$
2.7. Tope poset and generating function
Fix a base chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$. This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$. The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by rank function $\rho_{B}(C)=d(B,C)$. In terms of notations in Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
Proposition 2.34 (Theorem 2.2 of [10], see also Theorem 4.4.2 of [4]).
An open interval $(C_{1},C_{2})$ inside $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\mathrm{rank}\mathcal{A}-\dim F-2}$, if the closed interval $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$; contractible otherwise.
Corollary 2.35 (Theorem 1.11 of [11], see also Corollary 4.4.3 of [4]).
The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is
$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\mathrm{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F},\exists F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
Proposition 2.36 (Theorem 3.4 of [3]).
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A},B)$ is a lattice for any $B\in\mathrm{Ch}(\mathcal{A})$.
Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch(\mathcal{A})}}q^{d(B,C)}$ be the rank generation function of $\mathcal{T}(\mathcal{A},B)$. Note that it is the sum of entries of the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
Proposition 2.37 ([29]).
If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then
$D_{\mathcal{A},B}(q)=\prod_{i=1}^{\ell}[1+e_{i}]_{q}$
for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-analog.
The generating function decomposition formula is generalized to supersolvable arrangements [3], inductively factored arrangements [18]. See also [1, 26, 12].
3. Magnitude of tope graphs
For a real hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of $\mathcal{T}(\mathcal{A})$.
3.1. Basic properties
We first prove a reciprocity result.
Proposition 3.1.
Let $\mathcal{A}$ be a central hyperplane arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
$\mathrm{Mag}(\mathcal{A},q^{-1})=q^{N}\mathrm{Mag}(\mathcal{A},q).$
Proof.
Recall that there is an involution $\tau:\mathrm{Ch}(\mathcal{A})\to\mathrm{Ch}(\mathcal{A})$ sending $D$ to D-D. For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N. Then the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
$q^{N}\mathcal{V}_{q^{-1}}=[q^{N-d(C,D)}]=[q^{d(C,-D)}]=\mathcal{V}_{q}P_{\tau},$
where $P_{\tau}$ is the permutation matrix induced by $\tau$. By definition of magnitude,
$\mathrm{Mag}(\mathcal{A},q^{-1})=\mathbf{1}^{T}(\mathcal{V}_{q^{-1}})^{-1}\mathbf{1}=\mathbf{1}^{T}(q^{-N}\mathcal{V}_{q}P_{\tau})^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}P_{\tau}^{-1}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathrm{Mag}(\mathcal{A},q),$
where we used $\mathbf{1}^{T}P_{\tau}^{-1}=\mathbf{1}^{T}$. ∎
Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ has the one-point property by [22].
Proposition 3.2.
Let $\mathcal{A}$ be a central hyperplane arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
Next we prove some main structural results of $\mathrm{Mag}(\mathcal{A})$.
Theorem 3.3.
For central arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form, with the leading coefficient of $Q(q)$ positive. Then
  1. (1)
    $\mathcal{A}=\deg Q-\deg P$.
  2. (2)
    $P,Q$ are palindromic.
  3. (3)
    Roots of $Q$ are all roots of unity, excluding $1$.
Proof.
Let $P(q)=a_{m}q^{m}+\cdots+a_{0}$ and $Q(q)=b_{n}q^{n}+\cdots+b_{0}$ where $a_{m}\neq 0$ and $b_{n}\neq 0$. Then $m=\deg P$ and $n=\deg Q$. Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$. By the reciprocity formula (Proposition 3.1),
$q^{N}=\frac{\mathrm{Mag}(\mathcal{A},q^{-1})}{\mathrm{Mag}(\mathcal{A},q)}=\frac{P(q^{-1})Q(q)}{P(q)Q(q^{-1})}.$
Multiplying both sides by $q^{m-n}$, we have
$q^{N+m-n}=\frac{q^{m}P(q^{-1})Q(q)}{P(q)q^{n}Q(q^{-1})}=\frac{(a_{m}+\cdots+a_{0}q^{m})(b_{n}q^{n}+\cdots+b_{0})}{(a_{m}q^{m}+\cdots+a_{0})(b_{n}+\cdots+b_{0}q^{n})}.$
Comparing degrees, we conclude N+mn=0N+m-n=0 and (1) is proved.
Let us write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ as the reverse polynomials. We have proved
$\mathrm{Mag}(\mathcal{A},q)=\frac{P(q)}{Q(q)}=\frac{P^{*}(q)}{Q^{*}(q)}.$
Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, $P$ and $Q$ would share a common root $1/\alpha$, contradicting to the assumption $P/Q$ is reduced. Now we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$. Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
$Q^{*}(1)=Q(1)=rQ(1).$
Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, $Q(1)\neq 0$, we conclude $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, so (2) is proved.
Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, roots of $Q(q)$ are all roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.33. We have already seen $Q(1)\neq 0$. So (3) is proved. ∎
If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$. So by Proposition 2.5, we have a decomposition of magnitude.
Proposition 3.4.
$\mathrm{Mag}(\mathcal{A}_{1}\oplus\mathcal{A}_{2})=\mathrm{Mag}(\mathcal{A}_{1})\mathrm{Mag}(\mathcal{A}_{2})$.
Coxeter arrangements have the natural action of the Coxeter groups on the chambers, which makes the magnitude easy to compute.
Proposition 3.5.
For Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
$\mathrm{Mag}(\mathcal{A}_{W})=\frac{\#W}{\prod_{i}[1+e_{i}]_{q}}.$
Proof.
The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutohedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.29), hence is vertex-transitive. The result follows from Proposition 2.3 and Proposition 2.37. ∎
3.2. Face decomposition conjecture
For the purpose of reciprocity, let us define the following.
Definition 3.6.
We define the interior magnitude of $\mathcal{A}$ by
$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\mathrm{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
where the latter equality is Proposition 3.1.
Let us propose a main conjecture of the magnitude of arrangements.
Conjecture 3.7.
We have a face decomposition of the magnitude.
$\mathrm{Mag}(\mathcal{A})$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})$
$=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Equivalently,
$\mathrm{Mag}(\mathcal{A})$ $=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}$
$=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
Remark 3.8.
This formula is the discrete $q$-analogue of the decomposition of the Euler characteristic of a zonotope into its open faces. Conjecture 3.7 is verified by computer for all examples listed in Section 3.3 and the appendix, and by easy computations for Boolean arrangements and all rank $2$ arrangements.
Corollary 3.9.
Conjecture 3.7 implies that $\mathrm{Mag}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
3.3. Examples
We use the notion $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [14].
Example 3.10.
The near-pencil $\mathcal{A}(n,0)$ has tope graph $C_{2(n-1)}\Box K_{2}$ and hence
$\mathrm{Mag}(\mathcal{A}(n,0))=\frac{4(n-1)}{[2]_{q}^{2}[n-1]_{q}}.$
Example 3.11.
Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$, which can be computed from Proposition 3.5. In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
$\mathrm{Mag}(\mathcal{A}(6,1))$ $=\frac{24}{[4]_{q}!}=\frac{24}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)}$
$=24-72q+96q^{2}-72q^{3}+48q^{4}-72q^{5}+120q^{6}-144q^{7}+144q^{8}-144q^{9}+144q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(9,1))$ $=\frac{48}{[2]_{q}[4]_{q}[6]_{q}}=\frac{48}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)}$
$=48-144q+192q^{2}-192q^{3}+240q^{4}-336q^{5}+432q^{6}-528q^{7}+624q^{8}-720q^{9}+816q^{10}+\cdots$
More examples are in the appendix.
4. Magnitude homology of tope graphs
We continue denoting invariants of $\mathcal{T}(\mathcal{A})$ simply by invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}(\mathcal{A})$ and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$. Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
4.1. Small length
Let us start our study of $MH_{k,\ell}(\mathcal{A})$ from investigating small length $\ell\leq 2$.
Proposition 4.1.
Let $\mathcal{A}$ be a real central hyperplane arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
  1. (1)
    $MH_{0,0}(\mathcal{A})$ is free abelian of rank $\beta_{0,0}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A})$.
  2. (2)
    $MH_{1,1}(\mathcal{A})$ is free abelian of rank $\beta_{1,1}(\mathcal{A})=2\sum_{H\in\mathcal{A}}c^{H}$.
  3. (3)
    $MH_{2,2}(\mathcal{A})$ is free abelian of rank $\beta_{2,2}(\mathcal{A})=\beta_{1,1}(\mathcal{A})+4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$.
Proof.
(1) and (2) follow from (1) and (2) of Proposition 2.8 and noting that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.20 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$.
$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH_{2,\vec{a}}(\mathcal{A}).$
By Theorem 2.23, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets. For $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, where $a_{0},a_{1}$ are chambers at distance $2$, we may suppose $S(a_{0},a_{1})=\{H_{1},H_{2}\}$ and let $X=H_{1}\cap H_{2}\in L_{2}(\mathcal{A})$. If $\mathcal{A}_{X}=2$, then the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ consists of two incomparable chambers between $a_{0}$ and $a_{1}$. If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton. Since $MH_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology, only the first type $(a_{0},a_{1})$ with $\mathcal{A}_{X}=2$ contributes to the magnitude homology a copy of $\mathbb{Z}$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial. So we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent. The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty. Then $MH_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$. Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$. Now (3) is proved.
To finish the proof, we need to show other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.20 again
$MH_{1,2}(A)\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH_{1,\vec{a}}(\mathcal{A}).$
Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty. Hence $MH_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$. ∎
4.2. Geodesic magnitude homology and diagonal tope graphs
Let us determine the geodesic magnitude homology of arrangements.
Theorem 4.2.
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is as follow
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathrm{rank}\mathcal{A}_{F}=k,\#\mathcal{A}_{F}=\ell}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}$
Proof.
By Corollary 2.22,
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{a,b\in\mathrm{Ch}(\mathcal{A}):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
By Proposition 2.34, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\mathrm{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$ and contractible otherwise. Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\mathrm{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contribute to $MH_{k,\ell}^{geod}(\mathcal{A})$ a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ and there is no other summand. ∎
Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ and let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. We can give a combinatorial description of geodesic magnitude Betti number.
Corollary 4.3.
The geodesic magnitude Betti number
$\beta_{k,\ell}^{geod}(\mathcal{A})=\sum_{X\in L_{k}(\mathcal{A}):\#\mathcal{A}_{X}=\ell}c^{X}c_{X}.$
In particular, geodesic magnitude homology is determined by $L(\mathcal{A})$.
As an application, we give a characterization of diagonal tope graphs.
Corollary 4.4.
Let $\mathcal{A}$ be an essential arrangement. Then $\mathcal{T}(\mathcal{A})$ is diagonal if and only if $\mathcal{A}$ is Boolean.
Proof.
We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14). Suppose $\mathcal{A}$ is not Boolean, then $n=\#\mathcal{A}>r=\mathrm{rank}\mathcal{A}$. Therefore
$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{geod}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
where the equation is from Corollary 4.3. This means $\mathcal{A}$ is not diagonal. ∎
4.3. Magnitude homology with support
Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, let us define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$. Notice that $S(\vec{x})\leq\ell(\vec{x})$. It is not difficult to verify that $x_{i}$ is smooth in $\vec{x}$, i.e. $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, is equivalent to $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$. So if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$. In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes. This motivated us to define the following.
Definition 4.5.
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$. The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$. Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
Definition 4.6.
When $\mathcal{B}=\mathcal{A}$, we call $MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude complex of $\mathcal{A}$ and $MH_{*,\ell}^{\circ}(\mathcal{A}):=MH_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude homology of $\mathcal{A}$. Let us also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
The following result is a direct consequence of the above observation that the magnitude boundary operator preserves crossed hyperplanes.
Proposition 4.7.
For arrangement $\mathcal{A}$, we have the decomposition of magnitude chain groups
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}),$
inducing a decomposition of homology groups
$MH_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MH_{k,\ell}(\mathcal{A};\mathcal{B}).$
Furthermore $MC_{k,\ell}(\mathcal{A};\mathcal{B})=0$ and $MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
4.4. Face decomposition of magnitude homology
We would like to further decompose the magnitdue complex/homology with support. To this end, we introduce another notation.
Definition 4.8.
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$. Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$. We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$,
$P_{k,\ell}(G;\mathcal{B})=\{(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A};\mathcal{B})\mid x_{0},\ldots,x_{k}\in V(G)\}.$
Lemma 4.9.
For $F\in\mathcal{F}(\mathcal{A})$, the chain groups $MC_{*,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$ form a subcomplex of $MC_{*,\ell}(\mathcal{A};\mathcal{B})$.
Proof.
We must show that the boundary operator $\partial$ stabilizes the subgroup. By Proposition 4.7, $\partial$ strictly preserves the set $\mathcal{B}$ of crossed hyperplanes. Furthermore, by Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2). Because the boundary operator $\partial_{i}$ merely deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained within the vertex set of $\mathcal{T}(\mathcal{A})_{F}$. Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$. ∎
Now we are ready to state the main structural result of the magnitude homology of an arrangement.
Theorem 4.10.
For an arrangement $\mathcal{A}$, we have the face decomposition of its magnitude complex
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}),$
inducing the face decomposition of magnitude homology
$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
Proof.
By Proposition 4.7, the global magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}).$
For a fixed subarrangement $\mathcal{B}$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$. By Definition 4.5, $S(\vec{x})=\mathcal{B}$, meaning the chain never crosses any hyperplane in the complement $\mathcal{A}\setminus\mathcal{B}$. Therefore, every chamber in $\vec{x}$ shares a strictly constant, non-zero sign on all hyperplanes in $\mathcal{A}\setminus\mathcal{B}$. This constant sign pattern uniquely defines a face $F\in\mathcal{F}(\mathcal{A})$ such that $F_{H}=0$ for $H\in\mathcal{B}$, and $F_{H}\in\{+,-\}$ for $H\notin\mathcal{B}$. By definition, the subarrangement associated with this face is exactly $\mathcal{A}_{F}=\mathcal{B}$. Because the sign pattern on $\mathcal{A}\setminus\mathcal{B}$ is uniquely determined for each chain, the subgroup $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ further decomposes into a direct sum over the faces $F$ that satisfy $\mathcal{A}_{F}=\mathcal{B}$,
$MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathcal{A}_{F}=\mathcal{B}}MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}).$
By Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph isometrically isomorphic to the tope graph of the localized subarrangement $\mathcal{A}_{F}$. Under this isomorphism, the $\mathcal{B}$-supported complex maps identically to the full interior complex of $\mathcal{A}_{F}$. Therefore, by Definition 4.6,
$MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})\cong MC_{k,\ell}(\mathcal{A}_{F};\mathcal{A}_{F})=MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Substituting this isomorphism back into the original sum over subarrangements $\mathcal{B}$ yields the result
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}\left(\bigoplus_{F:\mathcal{A}_{F}=\mathcal{B}}MC^{\circ}_{k,\ell}(\mathcal{A}_{F})\right)=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Since this is a decomposition into subcomplexes, we have the face decomposition of magnitude homology
$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
For an arrangement $\mathcal{A}$, let us write
$\chi^{\circ}(\mathcal{A},q):=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}\in\mathbb{Z}[\![q]\!].$
Then we have a reformulation of Conjecture 3.7.
Corollary 4.11.
Let $r=\mathrm{rank}\mathcal{A}$ and $n=\#\mathcal{A}$. The following are equivalent
  1. (1)
    $\mathrm{Mag}(\mathcal{A},q)=\sum_{F}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q)$;
  2. (2)
    $\chi^{\circ}(\mathcal{A},q)=\mathrm{Mag}^{\circ}(\mathcal{A},q)$;
  3. (3)
    $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$.
Proof.
Since Euler characteristic is additive over direct sums, the face decomposition of magnitude homology (Theorem 4.10) yields
$\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi_{\ell}^{\circ}(\mathcal{A}_{F}).$
Multiplying by $q^{\ell}$ and summing over $\ell$, we obtain
$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}(\mathcal{A}_{F},q),$
By Möbius inversion on $\mathcal{F}(\mathcal{A})$, we see that (1) is equivalent to (2). The equivalence between (2) and (3) is straightforward by definition. ∎
Remark 4.12.
Note that if there is a (quasi-)isomorphism of chain complexes
$MC_{*,\ell}(\mathcal{A})\to MC_{*+r,\ell+n}^{\circ}(\mathcal{A})=MC_{*+r,\ell+n}(\mathcal{A};\mathcal{A})$
for all $\ell\geq 0$, then the conditions in Corollary 4.11 hold.
4.5. Magnitude homology on the diagonal
In this subsection, we give a combinatorial formula of the magnitude homology of arrangements on the diagonal $MH_{\ell,\ell}(\mathcal{A})$.
Using Theorem 4.10, we can determine the interior magnitude homology of Boolean arrangements.
Corollary 4.13.
The interior magnitude homology of Boolean arrangement $MH_{k,\ell}^{\circ}(Bl_{d})$ concentrates on the diagonal $k=\ell$ and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
$\beta_{\ell,\ell}^{\circ}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$.
Proof.
We know from Corollary 4.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$. By Theorem 4.10, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence concentrates on the diagonal $k=\ell$. Theorem 4.10 also gives
$\beta_{\ell,\ell}(Bl_{d})=\sum_{F\in\mathcal{F}(Bl_{d})}\beta_{\ell,\ell}^{\circ}((Bl_{d})_{F}).$
Notice that if $F\in\mathcal{F}(Bl_{d})$ has $\mathrm{codim}F=j$, the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$. Then applying Möbius inversion on $\mathcal{F}(\mathcal{A})$ yields the desired formula. ∎
Let us prove the following vanishing result of diagonal interior magnitude homology for non-Boolean arrangements.
Theorem 4.14.
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Suppose $\mathcal{A}$ is not Boolean $r<n$, then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes for all length
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\forall\ell\geq 0.$
Proof.
Choose an isometric embedding $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}=\{+,-\}^{n}$ of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{n}$, which is not surjective since $r<n$. Since $\phi$ is distance preserving, it induces injective chain maps $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{n})$. Let us take $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})\cap\ker\partial\subseteq MC_{\ell,\ell}(\mathcal{A})$. Since $\partial c=0$, $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{n})$ is a cycle. For fixed $\ell$, $MC_{\ell+1,\ell}(Q_{n})=0$, by abuse of notation, we say $\phi_{\#}(c)$ is a homology class in $MH_{\ell,\ell}(Q_{n})$. Recall the isomorphism given in Example 2.14
$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{n}),$
is induced by the Eilenberg–Zilber chain map [16]. We may assume there are basis elements $b^{\lambda}$ of the left hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$. We must show that $d_{\lambda}=0$.
Since $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})$, it is a linear combination of walks that cross all hyperplanes in $\mathcal{A}$. Its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{n})$ is a linear combination of walks in $Q_{n}$ along which every coordinate gets flipped at least once. This means the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$. We write such $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$. So $b^{\lambda}$ consists of totally $\ell$ steps in these $n$ dimensions with distribution $\lambda$.
By definition of the Eilenberg-Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps. It is a linear combination of walks in $Q_{n}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminating vertex in $Q_{1}$ respectively. Since all $\ell_{i}\geq 1$, any vertex in $Q_{n}$ must be visited by at least one walk in $\nabla(b^{\lambda})$. The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk due to distribution difference.
Now recall that $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}$ is not surjective, there is a vertex $v_{0}\in Q_{n}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$. We observed that any $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 2), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$. We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$. Therefore we conclude $c=0$, and $MH_{\ell,\ell}^{\circ}(\mathcal{A})=0$.
$G^{+}(+,+,+)$$G^{-}(-,-,-)$(+,,)(+,-,-)Start: (+,+,)(+,+,-)(,+,)(-,+,-)(,+,+)(-,+,+)(,,+)(-,-,+)End: (+,,+)(+,-,+)Step 1 ($Z_{1}$)Step 2 ($X_{1}$)Step 3 ($X_{2}$)Step 4 ($Y_{1}$)
Figure 2. The tope graph of the $A_{2}$ arrangement embedded as a $C_{6}$ hexagon (blue/bold) inside the Boolean $Q_{3}$ cube. Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$. A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown hitting the forbidden vertex $G^{+}$. Also notice that this walk cannot appear in other $\nabla(b^{\mu})$.
Remark 4.15.
By Corollary 4.13 and Theorem 4.14, we see that $\mathcal{A}$ is Boolean if and only if $\beta_{\ell,\ell}^{\circ}(\mathcal{A})\neq 0$ for some $\ell\geq 1$.
Corollary 4.16.
Let $\mathcal{A}$ be an arrangement. The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$. Precisely, for $\ell\geq 1$,
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\mathrm{rank}X-1}$.
Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$, where $d_{max}=\max\{\mathrm{rank}X\mid X\in L(\mathcal{A}),\mathrm{rank}X=\#\mathcal{A}_{X}\}$.
Proof.
Since $MC_{\ell+1,\ell}(\mathcal{A})=0$, the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian. By Theorem 4.10, the global magnitude homology admits a face decomposition. By restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F}).$
By Theorem 4.14, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e. $\mathcal{A}_{F}=\mathrm{rank}\mathcal{A}_{F}$. Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\mathrm{rank}X=\#\mathcal{A}_{X}$. For these faces, $\mathcal{A}_{F}\cong Bl_{\mathrm{rank}X}$. We can group the non-zero terms of the summation by their support flats $X\in L(\mathcal{A})$. For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$. Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 4.13), $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\mathrm{rank}X$, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
For $\ell\geq\mathrm{rank}X$, the binomial $\binom{\ell-1}{\mathrm{rank}X-1}$ is a polynomial in $\ell$ of degree $\mathrm{rank}X-1$. Therefore for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$. ∎
Remark 4.17.
Example 2.14 and Proposition 4.1 (2)(3) are special cases of the formula in Corollary 4.16.
5. Open problems
Besides the face decomposition conjecture (Conjecture 3.7, see also Corollary 4.11 and Remark 4.12), we propose the following conjectures based on observations of examples in the appendix.
Conjecture 5.1.
Let $\mathcal{A}$ be an arrangement of rank $r$ and $n=\#\mathcal{A}$. Let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form. Then
  1. (1)
    Suppose $r\geq 3$ is odd. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{2n}(q)$ divides $Q(q)$.
  2. (2)
    Suppose $r\geq 4$ is even. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{n}(q)$ divides $Q(q)$.
  3. (3)
    The coefficients of the power series expression of $\mathrm{Mag}(\mathcal{A},q)$ have alternating signs (equivalently $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0,\forall\ell\geq 0$).
  4. (4)
    $MH_{k,\ell}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
  5. (5)
    $MH_{k,\ell}(\mathcal{A})$ is torsion free.
  6. (6)
    If $\mathcal{A}$ is not Boolean, then $\beta_{0,0}(\mathcal{A})=\beta_{r,n}(\mathcal{A})$.
Acknowledgement
The author thanks Weili Guo and Masahiko Yoshinaga for comments on the preliminary draft of this paper.
Appendix A Computer experiments
The following examples are computed by using SageMath [28].
Example A.1.
We list the results of magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [14] for $n\leq 13$. Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.10 and 3.11. We use the database of [6] to pull the list of normal vectors.
$\mathrm{Mag}(\mathcal{A}(7,1))$ $=\frac{32q^{4}-40q^{2}+32}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{[2]_{q}^{2}[3]_{q}(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{14}(q)}$
$=32-96q+120q^{2}-72q^{3}+24q^{4}-48q^{5}+144q^{6}-272q^{7}+360q^{8}-336q^{9}+240q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(8,1))$ $=\frac{40q^{6}-8q^{4}-16q^{3}-8q^{2}+40}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{[4]_{q}!(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{16}(q)}$
$=40-120q+152q^{2}-112q^{3}+88q^{4}-136q^{5}+240q^{6}-344q^{7}+352q^{8}-248q^{9}+176q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,1))$ $=\frac{60q^{8}+60q^{7}-20q^{6}-40q^{5}-40q^{3}-20q^{2}+60q+60}{(q^{8}+4q^{7}+8q^{6}+11q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1)(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{[2]_{q}^{2}[3]_{q}[5]_{q}(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{20}(q)}$
$=60-180q+220q^{2}-140q^{3}+60q^{4}-80q^{5}+220q^{6}-380q^{7}+420q^{8}-340q^{9}+220q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,2))$ $=\mathrm{Mag}(\mathcal{A}(10,3))=\frac{60q^{6}-12q^{4}-48q^{3}-12q^{2}+60}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{[4]_{q}!(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)^{2}\Phi_{20}(q)}$
$=60-180q+228q^{2}-192q^{3}+204q^{4}-300q^{5}+432q^{6}-564q^{7}+660q^{8}-672q^{9}+576q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(11,1))$ $=\frac{72q^{6}-16q^{4}-64q^{3}-16q^{2}+72}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{[4]_{q}!(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{22}(q)}$
$=72-216q+272q^{2}-232q^{3}+256q^{4}-376q^{5}+528q^{6}-680q^{7}+800q^{8}-824q^{9}+784q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,1))$ $=\frac{84q^{6}-84q^{5}-36q^{4}+96q^{3}-36q^{2}-84q+84}{(q^{6}+2q^{5}+2q^{4}+2q^{3}+2q^{2}+2q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{[2]_{q}[6]_{q}(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{6}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+300q^{2}-168q^{3}+36q^{4}-84q^{5}+336q^{6}-588q^{7}+636q^{8}-504q^{9}+372q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,2))$ $=\frac{84q^{10}+84q^{9}+64q^{8}-4q^{7}-44q^{6}-128q^{5}-44q^{4}-4q^{3}+64q^{2}+84q+84}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{[5]_{q}!(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+316q^{2}-260q^{3}+252q^{4}-344q^{5}+508q^{6}-692q^{7}+804q^{8}-796q^{9}+760q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,3))$ $=\frac{84q^{6}-12q^{4}-96q^{3}-12q^{2}+84}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{[4]_{q}!(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+324q^{2}-312q^{3}+396q^{4}-564q^{5}+720q^{6}-876q^{7}+1044q^{8}-1128q^{9}+1116q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,1))$ $=\frac{96q^{8}-96q^{7}+72q^{6}-48q^{5}-48q^{3}+72q^{2}-96q+96}{(q^{8}+2q^{7}+3q^{6}+4q^{5}+4q^{4}+4q^{3}+3q^{2}+2q^{1}+1)(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{[2]_{q}^{2}(1+2q^{2}+2q^{4}+q^{6})(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)\Phi_{26}(q)}$
$=96-288q+360q^{2}-288q^{3}+264q^{4}-384q^{5}+624q^{6}-864q^{7}+984q^{8}-960q^{9}+888q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,2))$ $=\frac{96q^{6}-144q^{3}+96}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{[4]_{q}!(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=96-288q+384q^{2}-432q^{3}+624q^{4}-864q^{5}+1008q^{6}-1152q^{7}+1392q^{8}-1584q^{9}+1632q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,3))$ $=\frac{96q^{10}+96q^{9}+80q^{8}-8q^{7}-64q^{6}-160q^{5}-64q^{4}-8q^{3}+80q^{2}+96q+96}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{[5]_{q}!(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{26}(q)}$
$=96-288q+368q^{2}-328q^{3}+336q^{4}-424q^{5}+584q^{6}-784q^{7}+912q^{8}-920q^{9}+920q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,4))$ $=\frac{104q^{6}-56q^{4}-48q^{3}-56q^{2}+104}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{[4]_{q}!(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=104-312q+360q^{2}-192q^{3}+72q^{4}-168q^{5}+432q^{6}-696q^{7}+792q^{8}-672q^{9}+504q^{10}+\cdots$
Example A.2.
Consider the rank $3$ generic arrangement $\mathcal{A}=U_{3,4}$ defined by $xyz(x+y+z)$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{14q^{2}-20q+14}{(q^{2}+2q+1)(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{[2]_{q}^{2}(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{\Phi_{2}(q)^{2}\Phi_{8}(q)}$
$=14-48q+96q^{2}-144q^{3}+178q^{4}-192q^{5}+192q^{6}-192q^{7}+206q^{8}-240q^{9}+288q^{10}+\cdots$
Example A.3.
Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement $\mathcal{A}$ of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{18q^{4}-2q^{3}-8q^{2}-2q+18}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{[2]_{q}^{2}[3]_{q}(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{10}(q)}$
$=18-56q+88q^{2}-96q^{3}+104q^{4}-154q^{5}+248q^{6}-336q^{7}+376q^{8}-392q^{9}+450q^{10}+\cdots$
Example A.4.
Consider the graphic arrangement $\mathcal{A}$ of the graph $K_{5}\setminus e$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{[2]_{q}[4]_{q}[9]_{q}(q^{5}+1)}$
$=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{9}(q)\Phi_{10}(q)}$
$=96-396q+756q^{2}-924q^{3}+996q^{4}-1320q^{5}+1956q^{6}-2652q^{7}+3252q^{8}-3756q^{9}+4212q^{10}+\cdots$
where the numerator $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
Example A.5.
Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]),
$\mathrm{Mag}(\mathcal{A})$ $=\frac{6(17q^{8}-4q^{7}-8q^{6}-9q^{5}+44q^{4}-9q^{3}-8q^{2}-4q+17)}{\Phi_{2}(q)^{3}\Phi_{3}(q)^{2}\Phi_{9}(q)\Phi_{10}(q)}$
$=102-432q+864q^{2}-1176q^{3}+1584q^{4}-2628q^{5}+4344q^{6}-6048q^{7}+7200q^{8}-8130q^{9}+9684q^{10}+\cdots$
where the numerator $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
Example A.6.
We collect here some results of magnitude Betti numbers using SageMath codes of [16].
[…]
0 4 0 0 0 0 0 0 0 0
1 0 8 0 0 0 0 0 0 0
2 0 0 12 0 0 0 0 0 0
3 0 0 0 16 0 0 0 0 0
4 0 0 0 0 20 0 0 0 0
5 0 0 0 0 0 24 0 0 0
6 0 0 0 0 0 0 28 0 0
7 0 0 0 0 0 0 0 32 0
8 0 0 0 0 0 0 0 0 36
Table 1. Magnitude Betti numbers $\beta_{k,l}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
[…]
0 6 0 0 0 0 0 0 0 0
1 0 12 0 0 0 0 0 0 0
2 0 0 12 6 0 0 0 0 0
3 0 0 0 12 12 0 0 0 0
4 0 0 0 0 12 12 6 0 0
5 0 0 0 0 0 12 12 12 0
6 0 0 0 0 0 0 12 12 12
7 0 0 0 0 0 0 0 12 12
8 0 0 0 0 0 0 0 0 12
Table 2. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
[…]
0 8 0 0 0 0 0 0 0 0
1 0 16 0 0 0 0 0 0 0
2 0 0 16 0 8 0 0 0 0
3 0 0 0 16 0 16 0 0 0
4 0 0 0 0 16 0 16 0 8
5 0 0 0 0 0 16 0 16 0
6 0 0 0 0 0 0 16 0 16
7 0 0 0 0 0 0 0 16 0
8 0 0 0 0 0 0 0 0 16
Table 3. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
[…]
0 24 0 0 0 0 0 0 0 0
1 0 72 0 0 0 0 0 0 0
2 0 0 96 48 0 0 0 0 0
3 0 0 0 120 96 0 24 0 0
4 0 0 0 0 144 96 48 72 0
5 0 0 0 0 0 168 96 96 96
6 0 0 0 0 0 0 192 96 96
7 0 0 0 0 0 0 0 216 96
8 0 0 0 0 0 0 0 0 240
Table 4. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{3}$.
[…]
0 120 0 0 0 0 0 0
1 0 480 0 0 0 0 0
2 0 0 840 360 0 0 0
3 0 0 0 1200 960 0 240
4 0 0 0 0 1560 1440 360
5 0 0 0 0 0 1920 1920
6 0 0 0 0 0 0 2280
Table 5. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{4}$.
[…]
0 14 0 0 0 0 0 0 0 0
1 0 48 0 0 0 0 0 0 0
2 0 0 96 0 0 0 0 0 0
3 0 0 0 144 14 0 0 0 0
4 0 0 0 0 192 48 0 0 0
5 0 0 0 0 0 240 96 0 0
6 0 0 0 0 0 0 288 144 14
7 0 0 0 0 0 0 0 336 192
8 0 0 0 0 0 0 0 0 384
Table 6. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$ Example A.2.
[…]
0 30 0 0 0 0 0 0
1 0 140 0 0 0 0 0
2 0 0 380 0 0 0 0
3 0 0 0 780 0 0 0
4 0 0 0 0 1340 30 0
5 0 0 0 0 0 2060 140
6 0 0 0 0 0 0 2940
Table 7. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{4,5}$.
[…]
0 18 0 0 0 0 0 0 0 0
1 0 56 0 0 0 0 0 0 0
2 0 0 88 24 0 0 0 0 0
3 0 0 0 120 48 18 0 0 0
4 0 0 0 0 152 48 80 0 0
5 0 0 0 0 0 184 48 136 24
6 0 0 0 0 0 0 216 48 168
7 0 0 0 0 0 0 0 248 48
8 0 0 0 0 0 0 0 0 280
Table 8. Magnitude Betti numbers $\beta_{k,l}$ for the graphic arrangement of $K_{4}\setminus e$ Example A.3
[…]
0 32 0 0 0 0 0 0 0 0
1 0 96 0 0 0 0 0 0 0
2 0 0 120 72 0 0 0 0 0
3 0 0 0 144 144 0 0 32 0
4 0 0 0 0 168 144 72 0 96
5 0 0 0 0 0 192 144 144 0
6 0 0 0 0 0 0 216 144 144
7 0 0 0 0 0 0 0 240 144
8 0 0 0 0 0 0 0 0 264
Table 9. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(7,1)$
[…]
0 40 0 0 0 0 0 0 0 0
1 0 120 0 0 0 0 0 0 0
2 0 0 152 72 16 0 0 0 0
3 0 0 0 184 144 32 0 0 40
4 0 0 0 0 216 144 104 0 16
5 0 0 0 0 0 248 144 176 0
6 0 0 0 0 0 0 280 144 176
7 0 0 0 0 0 0 0 312 144
8 0 0 0 0 0 0 0 0 344
Table 10. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(8,1)$
[…]
0 48 0 0 0 0 0 0 0 0
1 0 144 0 0 0 0 0 0 0
2 0 0 192 48 48 0 0 0 0
3 0 0 0 240 96 96 0 0 0
4 0 0 0 0 288 96 144 0 48
5 0 0 0 0 0 336 96 192 0
6 0 0 0 0 0 0 384 96 192
7 0 0 0 0 0 0 0 432 96
8 0 0 0 0 0 0 0 0 480
Table 11. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(9,1)$
Reference
  • [1] T. Abe, T. Horiguchi, M. Masuda, S. Murai, and T. Sato (2020) Hessenberg varieties and hyperplane arrangements. J. Reine Angew. Math. 764, pp. 241–286. External Links: ISSN 0075-4102,1435-5345, Document, Link, MathReview (Marko Radovanović) Cited by: §2.7.
  • [2] M. Aguiar and S. Mahajan (2017) Topics in hyperplane arrangements. Mathematical Surveys and Monographs, Vol. 226, American Mathematical Society, Providence, RI. External Links: ISBN 978-1-4704-3711-4, Document, Link, MathReview (Tom C. Braden) Cited by: §2.6, Proposition 2.2, Remark 2.25, Theorem 2.33.
  • [3] A. Björner, P. H. Edelman, and G. M. Ziegler (1990) Hyperplane arrangements with a lattice of regions. Discrete Comput. Geom. 5 (3), pp. 263–288. External Links: ISSN 0179-5376,1432-0444, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7, Proposition 2.36.
  • [4] A. Björner, M. Las Vergnas, B. Sturmfels, N. White, and G. M. Ziegler (1999) Oriented matroids. Second edition, Encyclopedia of Mathematics and its Applications, Vol. 46, Cambridge University Press, Cambridge. External Links: ISBN 0-521-77750-X, Document, Link, MathReview Entry Cited by: Theorem 1.1, Proposition 2.30, Proposition 2.34, Corollary 2.35.
  • [5] S. Brenner, J. Cardinal, T. McConville, A. Merino, and T. Mütze (2026) Combinatorial generation via permutation languages. VII. Supersolvable hyperplane arrangements. European J. Combin. 135, pp. Paper No. 104367. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview Entry Cited by: §2.5.
  • [6] M. Cuntz, S. Elia, and J. Labbé (2022) Congruence normality of simplicial hyperplane arrangements via oriented matroids. Ann. Comb. 26 (1), pp. 1–85. External Links: ISSN 0218-0006,0219-3094, Document, Link, MathReview (Hery Randriamaro) Cited by: Example A.1.
  • [7] G. Denham, G. Dorpalen-Barry, and N. Proudfoot (2026) Cleanliness and the varchenko-gelfand algebra. arXiv preprint arXiv:2512.10077v2. Cited by: Example A.5.
  • [8] G. Denham and P. Hanlon (1999) Some algebraic properties of the Schechtman-Varchenko bilinear forms. In New perspectives in algebraic combinatorics (Berkeley, CA, 1996–97), Math. Sci. Res. Inst. Publ., Vol. 38, pp. 149–176. External Links: ISBN 0-521-77087-4, MathReview (P. Orlik) Cited by: §2.6.
  • [9] R. Diestel ([2025] ©2025) Graph theory. Sixth edition, Graduate Texts in Mathematics, Vol. 173, Springer, Berlin. External Links: ISBN 978-3-662-70106-5; 978-3-662-70107-2, MathReview Entry Cited by: §2.5.
  • [10] P. H. Edelman and J. W. Walker (1985) The homotopy type of hyperplane posets. Proc. Amer. Math. Soc. 94 (2), pp. 221–225. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Udo Pachner) Cited by: Proposition 2.34.
  • [11] P. H. Edelman (1984) A partial order on the regions of ${\bf R}^{n}$ dissected by hyperplanes. Trans. Amer. Math. Soc. 283 (2), pp. 617–631. External Links: ISSN 0002-9947,1088-6850, Document, Link, MathReview (Robert A. Proctor) Cited by: Corollary 2.35.
  • [12] L. Giordani, T. Möller, P. Mücksch, and G. Roehrle (2025) On connected subgraph arrangements. arXiv preprint arXiv:2502.18144v1. Cited by: §2.7.
  • [13] K. Gomi (2025) Magnitude homology of geodesic space. arXiv preprint arXiv:1902.07044v3. Cited by: Proposition 2.21.
  • [14] B. Grünbaum (2009) A catalogue of simplicial arrangements in the real projective plane. Ars Math. Contemp. 2 (1), pp. 1–25. External Links: ISSN 1855-3966,1855-3974, Document, Link, MathReview Entry Cited by: Example A.1, §1, §3.3.
  • [15] Y. Gu (2018) Graph magnitude homology via algebraic morse theory. arXiv preprint arXiv:1809.07240v1. Cited by: Example 2.16.
  • [16] R. Hepworth and S. Willerton (2017) Categorifying the magnitude of a graph. Homology Homotopy Appl. 19 (2), pp. 31–60. External Links: ISSN 1532-0073,1532-0081, Document, Link, MathReview (Haimiao Chen) Cited by: Example A.6, §1, §2.3, Proposition 2.10, Theorem 2.12, Theorem 2.12, Proposition 2.8, Proposition 2.9, §4.5.
  • [17] W. Hochstättler and V. Welker (2019) The Varchenko determinant for oriented matroids. Math. Z. 293 (3-4), pp. 1415–1430. External Links: ISSN 0025-5874,1432-1823, Document, Link, MathReview (Daryl Funk) Cited by: §2.6.
  • [18] M. Jambu and L. Paris (1995) Combinatorics of inductively factored arrangements. European J. Combin. 16 (3), pp. 267–292. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7.
  • [19] R. Kaneta and M. Yoshinaga (2021) Magnitude homology of metric spaces and order complexes. Bull. Lond. Math. Soc. 53 (3), pp. 893–905. External Links: ISSN 0024-6093,1469-2120, Document, Link, MathReview (David Matthew Freeman) Cited by: §2.3, §2.3, Theorem 2.20, Theorem 2.23.
  • [20] V. Körber, T. Schnieders, J. Stricker, and J. Walizadeh (2025) Hamiltonian cycles in simplicial and supersolvable hyperplane arrangements. arXiv preprint arXiv:2508.14538v2. Cited by: §2.5.
  • [21] T. Leinster and M. W. Meckes (2017) The magnitude of a metric space: from category theory to geometric measure theory. In Measure theory in non-smooth spaces, Partial Differ. Equ. Meas. Theory, pp. 156–193. External Links: ISBN 978-3-11-055082-5; 978-3-11-055083-2, MathReview (Tobias W. Fritz) Cited by: §1.
  • [22] T. Leinster and M. Meckes (2023) Spaces of extremal magnitude. Proc. Amer. Math. Soc. 151 (9), pp. 3967–3973. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Shlomo Reisner) Cited by: §3.1.
  • [23] T. Leinster and M. Shulman (2021) Magnitude homology of enriched categories and metric spaces. Algebr. Geom. Topol. 21 (5), pp. 2175–2221. External Links: ISSN 1472-2747,1472-2739, Document, Link, MathReview (Ahmet A. Husainov) Cited by: §1.
  • [24] T. Leinster (2013) The magnitude of metric spaces. Doc. Math. 18, pp. 857–905. External Links: ISSN 1431-0635,1431-0643, MathReview (J. Böhm) Cited by: §1.
  • [25] T. Leinster (2019) The magnitude of a graph. Math. Proc. Cambridge Philos. Soc. 166 (2), pp. 247–264. External Links: ISSN 0305-0041,1469-8064, Document, Link, MathReview (Lorenzo Traldi) Cited by: §1, §2.2, Proposition 2.3, Proposition 2.5, Theorem 2.6, Proposition 2.7.
  • [26] S. Oh, A. Postnikov, and H. Yoo (2008) Bruhat order, smooth Schubert varieties, and hyperplane arrangements. J. Combin. Theory Ser. A 115 (7), pp. 1156–1166. External Links: ISSN 0097-3165,1096-0899, Document, Link, MathReview (David E. Anderson) Cited by: §2.7.
  • [27] P. Orlik and H. Terao (1992) Arrangements of hyperplanes. Grundlehren der mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], Vol. 300, Springer-Verlag, Berlin. External Links: ISBN 3-540-55259-6, Document, Link, MathReview (Michel Yves Jambu) Cited by: §2.4.
  • [28] The Sage Developers (2025) Sagemath, the Sage Mathematics Software System (Version 10.8). Note: https://www.sagemath.org Cited by: Appendix A.
  • [29] L. Solomon (1966) The orders of the finite Chevalley groups. J. Algebra 3, pp. 376–393. External Links: ISSN 0021-8693, Document, Link, MathReview (T. Ono) Cited by: Proposition 2.37.
  • [30] R. P. Stanley (2007) An introduction to hyperplane arrangements. In Geometric combinatorics, IAS/Park City Math. Ser., Vol. 13, pp. 389–496. External Links: ISBN 978-0-8218-3736-8; 0-8218-3736-2, Document, Link, MathReview Entry Cited by: §2.4, Remark 2.25.
  • [31] A. Varchenko (1993) Bilinear form of real configuration of hyperplanes. Adv. Math. 97 (1), pp. 110–144. External Links: ISSN 0001-8708,1090-2082, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.6, Theorem 2.33.
  • [32] Y. Yagi and M. Yoshinaga (2026) Reconstrucion of oriented matroids from varchenko-gelfand algebras. arXiv preprint arXiv:2509.19905v3. Cited by: §1.
  • [33] T. Zaslavsky (1975) Facing up to arrangements: face-count formulas for partitions of space by hyperplanes. Mem. Amer. Math. Soc. 1, pp. vii+102. External Links: ISSN 0065-9266,1947-6221, Document, Link, MathReview (E. Jucovič) Cited by: §2.4, Remark 2.25.