### 01 May 2026 v3

# Magnitude homology of real hyperplane arrangements

Abstract.
We define and study the magnitude and magnitude homology of a real hyperplane arrangement by regarding its tope graph as a metric space. We prove several structural results for the magnitude of arrangements, including a symmetry formula, palindromicity of the numerator and denominator, a face decomposition formula, and results on the sign pattern of the magnitude power series. For the magnitude homology of arrangements, we obtain combinatorial formulas for small lengths and show that it detects Boolean arrangements. We also lift the face decomposition formula to a homological decomposition and derive explicit formulas for the diagonal magnitude Betti numbers. Another notable feature is that the magnitude Euler characteristic satisfies a reciprocity theorem analogous to Ehrhart–Macdonald reciprocity. We conclude by presenting several conjectures. In particular, we conjecture that the magnitude homology of an arrangement is torsion-free and is determined by the intersection lattice.
Key words and phrases:
hyperplane arrangement, tope graph, magnitude, magnitude homology
2020 Mathematics Subject Classification:
Primary 52C35; Secondary 55N35
1. Introduction
Let $\mathcal{A}$ be an arrangement of hyperplanes passing through the origin in $\mathbb{R}^{d}$. The hyperplanes of $\mathcal{A}$ divide $\mathbb{R}^{d}$ into regions called chambers. We denote the set of chambers by $\mathrm{Ch}(\mathcal{A})$. The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ and with edges joining two chambers separated by a single hyperplane. More generally, the tope graph is defined for an oriented matroid. Although its definition is elementary, the tope graph contains essentially all combinatorial information about a hyperplane arrangement:
Theorem 1.1 (Theorem 4.2.14 of [4]).
A simple oriented matroid is uniquely determined (up to reorientation) by its unlabeled tope graph.
We refer to Figure 1 of [34] for a detailed diagram of implications. In particular, if two arrangements have isomorphic tope graphs, then they have homeomorphic complexified complements, isomorphic intersection lattices, and isomorphic filtered Varchenko–Gelfand algebras. Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
In this paper, by applying the emerging theory of magnitude to tope graphs, we define the magnitude and magnitude homology of arrangements and initiate the study of these invariants. Magnitude is a cardinality-like invariant of enriched categories and metric spaces, defined by Leinster [24, 21]; for metric spaces, it can be interpreted as measuring their “effective size.” The magnitude of a graph is defined by regarding the graph as a metric space via the shortest-path metric [25]. The magnitude of a graph is a rational function defined using the inverse of the $q$-distance matrix; in the case of hyperplane arrangements, this matrix coincides with the classical Varchenko matrix. The notion of magnitude homology, a categorification of magnitude, was introduced for graphs by Hepworth and Willerton [17], and for enriched categories and metric spaces by Leinster and Shulman [23]. It is a bigraded homology theory that reflects deep metric properties.
Our motivation is twofold. First, magnitude and magnitude homology give new invariants of real hyperplane arrangements. What information about an arrangement is reflected by these new invariants remains mysterious and is worth exploring. Second, the tope graphs of arrangements form a class of graphs with many distinctive features, such as antipodal symmetry and decompositions into faces, and may therefore serve as a source of interesting examples in the magnitude theory of graphs.
At the decategorified level, we show that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, satisfies remarkable structural properties. We prove a symmetry relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, which reflects the antipodal symmetry of the tope graph. The reduced rational form of $\mathrm{Mag}(\mathcal{A},q)$ has palindromic numerator and denominator, and all its poles are roots of unity. Moreover, if we define the interior magnitude by $\mathrm{Mag}^{\circ}(\mathcal{A},q)=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q)$, then the following face decomposition formula holds:
$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
Here $\mathcal{F}(\mathcal{A})$ denotes the set of faces (covectors) of $\mathcal{A}$, and $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subset H\}$. This serves as a discrete $q$-analogue of the classical decomposition of the Euler characteristic of a zonotope. We then investigate the sign pattern of the magnitude power series. In particular, we show that for arrangements of rank at most $3$, the signs of the magnitude coefficients eventually alternate. We also present an example with non-alternating magnitude.
The magnitude homology $MH_{k,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is a more refined invariant. By analyzing the structure of gated subgraphs arising from the faces of an arrangement, we establish the basic properties of this invariant. First, we describe all magnitude homology groups of length $\ell\leq 2$ in terms of the combinatorics of the arrangement. Next, we show that magnitude homology detects Boolean arrangements: an arrangement is Boolean if and only if its magnitude homology is concentrated on the diagonal. Furthermore, we define the interior magnitude homology $MH^{\circ}_{k,\ell}(\mathcal{A})$, which is a categorification of the interior magnitude $\mathrm{Mag}^{\circ}(\mathcal{A})$, and lift the face decomposition formula for magnitude to the homological level. In particular, we obtain the Euler characteristic reciprocity
$\chi^{\circ}_{\ell+\#\mathcal{A}}(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}}\chi_{\ell}(\mathcal{A}),$
where $\chi_{\ell}(\mathcal{A})$ (resp. $\chi^{\circ}_{\ell}(\mathcal{A})$) is the Euler characteristic of $MH_{*,\ell}(\mathcal{A})$ (resp. $MH^{\circ}_{*,\ell}(\mathcal{A})$). This can be regarded as a new combinatorial reciprocity theorem analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31]. As an application of the homological face decomposition, we give a complete description of the diagonal magnitude Betti numbers in terms of the intersection lattice.
The paper is organized as follows. Section 2 recalls the necessary material from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition. Section 3 investigates the basic properties and structure of the magnitude rational function. Section 4 focuses on the sign pattern of magnitude power series. Section 5 explores magnitude homology, presenting the Boolean diagonality theorem and the face decomposition theorem. Finally, Section 6 poses several open problems and conjectures for future study. The appendix contains results from numerous SageMath computer experiments.
Use of AI
In writing this paper, we used AI in the following ways.
  • A substantial part of the proof strategy for Theorem 3.7 was generated by ChatGPT-5.4 Pro.
  • Part of the proof strategy for Theorem 4.5 was generated by ChatGPT-5.5 Pro.
  • We used ChatGPT-5.5 Pro to proofread the manuscript for minor errors.
We emphasize, however, that we have not used any AI-generated text without careful verification. The text of this paper has been written under the full responsibility of the authors.
Convention
In this paper, all graphs considered are finite, simple, and undirected unless otherwise specified. When we discuss a topological property or construction of a poset, we mean that of its order complex. We shall use the term “arrangement” to mean a real central hyperplane arrangement.
2. Recollections
2.1. Metric graph theory
For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from a vertex $x\in V$ to a vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$, and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$. We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ to be the minimum length among walks from $x$ to $y$. If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$. Then $(G,d)$ is a generalized metric space. In particular, when $G$ is connected, that is, when there is a walk between any pair of vertices, $(G,d)$ is a metric space. We call the maximum value of $d$ the diameter of $G$, denoted by $\operatorname{diam}(G)$.
A subgraph $H$ of $G$ can be equipped with a metric in two ways. One is the restriction of $d_{G}$ to $H$, and the other is the graph metric $d_{H}$ on $H$ itself. We say that $H$ is an isometric subgraph of $G$ if the two metrics coincide, that is, if $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk. For $x,y\in V$, define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$. We equip it with the partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ that traverses $u$ before $v$. Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$. A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$. More generally, vertices $u_{1},\ldots,u_{k}$ form a strictly linearly ordered chain in $x,y]_{G}$, that is, $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$, if and only if
(2.1) $d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ obtained by excluding $x$ and $y$.
We say that a subgraph $H$ of $G$ is convex in $G$ if, for any $x,y\in V(H)$, every minimal walk from $x$ to $y$ traverses only vertices in $H$; equivalently, $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$. Every convex subgraph is an isometric subgraph, but the converse is false.
Definition 2.1.
Let $G$ be a connected graph, let $H$ be an isometric subgraph of $G$, and let $x\in V(G)$. We say that $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that, for any $v\in V(H)$,
$d(x,v)=d(x,u)+d(u,v).$
Such a vertex $u$ must be unique, since the above equation says that $u$ is the closest vertex to $x$ in $H$. We call $u$ the gate of $H$ with respect to $x$. Intuitively, every minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate. We say that $H$ is a gated subgraph of $G$ if it is gated with respect to every vertex of $G$. In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection. Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
Proposition 2.2 ([2, Proposition A.1]).
A gated subgraph is necessarily convex.
Proof.
Suppose $H$ is gated in $G$. Choose a minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$, and let $z$ be an arbitrary vertex in this walk. We need to show that $z\in V(H)$. Since $H$ is gated, we have
d(z,x)=d(z,p(z))+d(p(z),x),d(z,y)=d(z,p(z))+d(p(z),y).d(z,x)=d(z,p(z))+d(p(z),x),~d(z,y)=d(z,p(z))+d(p(z),y).
Since $z$ belongs to $x,y]_{G}$, adding these two equations yields
$d(x,y)=d(x,z)+d(z,y)=d(x,p(z))+d(p(z),y)+2d(z,p(z))\geq d(x,y)+2d(z,p(z)).$
This forces $d(z,p(z))=0$, i.e. $z=p(z)\in V(H)$. ∎
In this paper, the graph homomorphisms we use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, that is, maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$. Note that gate projections are distance non-increasing.
2.2. Graph magnitude
See Leinster [25] for details. For a finite simple graph $G=(V,E)$, define the Zeta matrix $Z_{G}=Z_{G}(q)=[q^{d(x,y)}]_{x,y\in V}$, where $q^{\infty}=0$ by convention. Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$. The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ can then be expressed as a power series in $\mathbb{Z}[\![q]\!]$. In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
The magnitude of $G$ is defined by the sum of all entries of $Z_{G}(q)^{-1}$.
$\mathrm{Mag}(G)=\mathrm{Mag}(G,q):=\sum_{x,y\in V}Z_{G}(q)^{-1}_{x,y}\in\mathbb{Q}(q)~(\mathrm{or}~\mathbb{Z}[\![q]\!]).$
In general, the sum of all entries of a matrix $A$ is equal to $\mathbf{1}^{T}A\mathbf{1}$, where $\mathbf{1}$ is the column vector of appropriate size with all entries equal to $1$. If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$. For a graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$. In other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$. Then the magnitude is given by
$\mathrm{Mag}(G)=\mathbf{1}^{T}w_{G}=\sum_{x\in V}w_{G}(x).$
The magnitude of a vertex-transitive graph can be computed explicitly.
Proposition 2.3 ([25, Lemma 3.2]).
If $G$ is vertex-transitive (i.e. $\mathrm{Aut}(G)$ acts transitively on $V$), then
$\mathrm{Mag}(G)=\frac{\#V}{\sum_{x\in V}q^{d(g,x)}}$
for any $g\in V$.
Proof.
The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and we denote it by $s=s(g)$. It is then easy to see that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$. We conclude $\mathrm{Mag}(G)=\#V/s$. ∎
Example 2.4.
Let $K_{n}$ be the complete graph on $n$ vertices and let $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Their magnitudes are given by
$\mathrm{Mag}(K_{n})$ $=\frac{n}{1+(n-1)q},$
$\mathrm{Mag}(C_{2n})$ $=\frac{2n(q-1)}{(q^{n}-1)(q+1)}=\frac{2n}{(1+q)(1+q+q^{2}+\cdots+q^{n-1})},$
$\mathrm{Mag}(C_{2n-1})$ $=\frac{(2n-1)(q-1)}{2q^{n}-q-1}=\frac{2n-1}{1+2q+2q^{2}+\cdots+2q^{n-1}}.$
Magnitude behaves like cardinality in the following sense.
Proposition 2.5 ([25, Lemmas 3.5 and 3.6]).
Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ for the disjoint union and Cartesian product of $G$ and $H$, respectively. Then
$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),\quad\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
Under certain conditions, magnitude satisfies inclusion-exclusion.
Theorem 2.6 ([25, Theorem 4.9]).
If a graph $X$ has a decomposition into subgraphs $X=G\cup H$ such that $G\cap H$ is convex in $X$ and gated in $H$, then
$\mathrm{Mag}(X)=\mathrm{Mag}(G)+\mathrm{Mag}(H)-\mathrm{Mag}(G\cap H).$
2.3. Magnitude homology of graphs
For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$. We say that a $k$-chain $\vec{x}$ is proper if adjacent nodes are distinct, that is, if $x_{i}\neq x_{i+1}$ for all $i$. Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and let $P_{k,\ell}(G)$ be the set of proper $k$-chains of length $\ell$. The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [17].
Proposition 2.7 ([25, Proposition 3.9]).
For a graph $G$, we have
$\mathrm{Mag}(G)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{\vec{x}\in P_{k}(G)}q^{\ell(\vec{x})}\in\mathbb{Z}[\![q]\!].$
In other words, if $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$, then
$c_{\ell}=\sum_{k=0}^{\ell}(-1)^{k}\#P_{k,\ell}(G).$
In particular, we have $c_{0}=\#V$ and $c_{1}=-2\#E$.
Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$. We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ by $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
$\partial_{i}(x_{0},\ldots,x_{k})=\begin{cases}(x_{0},\ldots,\widehat{x_{i}},\ldots,x_{k}),&\text{ if }d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1});\\ 0,&\text{ otherwise},\end{cases}$
where $\widehat{x_{i}}$ denotes deletion of $x_{i}$. One can verify that $\partial^{2}=0$. We call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$. Since the boundary map preserves endpoints, we have a decomposition
$MC_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MC_{k,\ell}(G;u,v),$
where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,\ x_{k}=v\}$. These subgroups form subcomplexes and induce a decomposition of homology groups
(2.2) $MH_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MH_{k,\ell}(G;u,v).$
We define the magnitude Betti number by $\beta_{k,\ell}(G)=\operatorname{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic by $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
Proposition 2.8 ([17, Propositions 2.9 and 2.10]).
Let $G=(V,E)$ be a graph.
  1. (1)
    $MH_{0,0}(G)$ is free abelian of rank $\beta_{0,0}(G)=\#V$.
  2. (2)
    $MH_{1,1}(G)$ is free abelian of rank $\beta_{1,1}(G)=2\#E$.
  3. (3)
    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\operatorname{diam}(G)\leq k\leq\ell$, and the inequality $\ell/\operatorname{diam}(G)\leq k$ is strict if $\operatorname{diam}(G)>1$ and $\ell>0$.
The following proposition is a straightforward consequence of Proposition 2.7 and says that magnitude can be recovered from magnitude homology.
Proposition 2.9 ([17, Theorem 2.8]).
Let $G$ be a graph, and $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$. Then
$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\operatorname{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
Hepworth–Willerton proved that magnitude homology is a functor.
Proposition 2.10 ([17, Proposition 3.3]).
Magnitude homology is a functor from the category of finite graphs with distance non-increasing maps to the category of bigraded abelian groups.
A direct application of functoriality is as follows. Let $G$ be a connected graph and let $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$. Let $\iota:H\to G$ be the inclusion. Both $p$ and $\iota$ are distance non-increasing. The composition $p\circ\iota=\text{id}_{H}$ induces $\text{id}=p_{*}\circ\iota_{*}:MH_{k,\ell}(H)\to MH_{k,\ell}(G)\to MH_{k,\ell}(H)$. This gives a decomposition of $MH_{k,\ell}(G)$.
Proposition 2.11.
If $p:G\to H$ is a gate projection, then
$MH_{k,\ell}(G)\cong MH_{k,\ell}(H)\oplus\ker(p_{*}).$
There is a Künneth formula for magnitude homology of graphs.
Theorem 2.12 ([17, Theorem 5.3]).
There is a natural short exact sequence that splits non-naturally
$0\to MH_{*,*}(G)\otimes MH_{*,*}(H)\xrightarrow{\square}MH_{*,*}(G\square H)\to\mathrm{Tor}(MH_{*-1,*}(G),MH_{*,*}(H))\to 0,$
where $\square$ is the exterior product defined in [17, Definition 5.2].
Example 2.13.
For the complete graph $K_{n}$, we have
$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=\ell;\\ 0,&\text{ otherwise}.\end{cases}$
The magnitude Betti number is given by $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify that $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
Example 2.14.
For the hypercube graph $Q_{d}=K_{2}^{\square d}$, the Künneth formula (Theorem 2.12) gives
$MH_{k,\ell}(Q_{d})=\bigoplus_{\sum k_{i}=k,\sum\ell_{i}=\ell}MH_{k_{1},\ell_{1}}(K_{2})\otimes\cdots\otimes MH_{k_{d},\ell_{d}}(K_{2}),$
Using the result for $K_{2}$, we obtain
$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=\ell;\\ 0,&\text{ otherwise},\end{cases}$
where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{\ell+d-1}{d-1}$. One can verify that $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
Definition 2.15.
We say that $G$ is diagonal if $MH_{k,\ell}(G)=0$ for $k\neq\ell$.
Example 2.16.
We have seen in Examples 2.13 and 2.14 that $K_{n}$ and $Q_{d}$ are diagonal. In particular, the cycle graph $C_{4}=Q_{2}$ is diagonal. However, Gu [16] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that the graphs $C_{n}$ ($n\geq 5$) are not diagonal.
Kaneta–Yoshinaga [20] constructed a computationally tractable approximation of magnitude homology. Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$. We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$. Every subchain of a geodesic proper chain is again geodesic and proper. Any proper $1$-chain is geodesic. A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$. If $u,v\in V(G)$ are vertices and $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
Definition 2.17.
We define the length $\ell$ geodesic magnitude complex by
$MC_{*,\ell}^{\mathrm{geod}}(G):=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MC_{*,\ell}(G;u,v).$
Its homology is called the length $\ell$ geodesic magnitude homology. The geodesic magnitude homology $MH^{\mathrm{geod}}_{k,\ell}(G)$ is a direct summand of $MH_{k,\ell}(G)$. It admits a decomposition
$MH_{k,\ell}^{\mathrm{geod}}(G)=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MH_{k,\ell}(G;u,v).$
Note that $MC_{*,\ell}^{\mathrm{geod}}(G)$ and $MH_{k,\ell}^{\mathrm{geod}}(G)$ are trivial if $\ell>\operatorname{diam}(G)$.
We define the geodesic magnitude Betti number by $\beta_{k,\ell}^{\mathrm{geod}}(G)=\operatorname{rank}MH_{k,\ell}^{\mathrm{geod}}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic by $\chi_{\ell}^{\mathrm{geod}}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\mathrm{geod}}(G)$. Observe that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{\mathrm{geod}}(G)$.
Geodesic magnitude homology is a coarse approximation of magnitude homology. As a more refined approximation, Kaneta–Yoshinaga [20] introduced geodesically simple magnitude homology. For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}$ ($i=1,2,\ldots,k-1$) is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$. A node that is not smooth is called a singular node. Note that endpoints are singular. We write $\varphi(\vec{x})$ for the subchain of $\vec{x}$ consisting of all singular nodes and call it the frame of $\vec{x}$. Note that $\varphi(\vec{x})$ is not necessarily proper.
Definition 2.18.
A proper chain $\vec{x}$ is geodesically simple if $\ell(\varphi(\vec{x}))=\ell(\vec{x})$.
Definition 2.19.
We define $MC_{k,\ell}^{\mathrm{gs}}(G)$ to be the subgroup of $MC_{k,\ell}(G)$ generated by geodesically simple chains. Then $MC_{*,\ell}^{\mathrm{gs}}(G)$ forms a subcomplex of $MC_{k,\ell}(G)$ [20, Propositions 3.6 and 3.7]. Its homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{\mathrm{gs}}(G)$.
How well geodesically simple magnitude homology approximates magnitude homology is controlled by the notion of a $4$-cut.
Definition 2.20.
A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic. Let $m_{G}$ be the minimum length of $4$-cuts of $G$. We set $m_{G}=\infty$ if $G$ has no $4$-cuts. Note that $m_{G}\geq 3$. Any proper chain with length smaller than $m_{G}$ is geodesically simple.
Lemma 2.21.
If $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$ satisfies $\varphi(\vec{x})=(x_{0},x_{k})$ and $\vec{x}$ is not geodesically simple, i.e. $(x_{i-1},x_{i},x_{i+1})$ is geodesic for $i=1,\ldots,k-1$, but $\vec{x}$ is not geodesic, then $\vec{x}$ has a subchain that is a $4$-cut.
Proof.
Let $\vec{y}=(y_{0},\ldots,y_{p})$ be a contiguous subchain of $\vec{x}$, that is, $y_{0}=x_{i},y_{1}=x_{i+1},\ldots,y_{p}=x_{i+p}$ for some $i$, such that $\vec{y}$ is not geodesic and $p$ is as small as possible. This means $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic. Then $p\geq 3$. We claim $\vec{z}:=(y_{0},y_{1},y_{p-1},y_{p})$ is a $4$-cut. First note that $y_{p-1}\neq y_{1}$ as they are connected by a geodesic $(y_{1},y_{2},\ldots,y_{p-1})$. That $(y_{0},y_{1},y_{p-1})$ and $(y_{1},y_{p-1},y_{p})$ are geodesic follows from the assumption that $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic. Moreover, $\ell(\vec{z})=d(y_{0},y_{1})+d(y_{1},y_{p-1})+d(y_{p-1},y_{p})=\sum_{i=1}^{p}d(y_{i-1},y_{i})>d(y_{0},y_{p})$, since $\vec{y}$ is not geodesic. This shows $\vec{z}$ is not geodesic and hence is a $4$-cut. ∎
Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$. We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$. Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and is empty if $\varphi(\vec{a})\neq\vec{a}$. We further define $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{\mathrm{gs}}(G)$. These groups form a subcomplex $MC^{\mathrm{gs}}_{*,\vec{a}}(G)$ whose homology $MH^{\mathrm{gs}}_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$. Kaneta–Yoshinaga proved that the chain complex $MC_{*,\ell}^{\mathrm{gs}}(G)$ decomposes as
(2.3) $MC_{k,\ell}^{\mathrm{gs}}(G)=\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MC^{\mathrm{gs}}_{k,\vec{a}}(G),$
and the same is true for the homology groups.
Theorem 2.22 ([20, Theorem 3.12]).
Let $G$ be a graph and $k,\ell>0$. We have a decomposition
$MH_{k,\ell}^{\mathrm{gs}}(G)\cong\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MH^{\mathrm{gs}}_{k,\vec{a}}(G).$
Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{\mathrm{gs}}(G)\cong MH_{k,\ell}(G)$.
Kaneta–Yoshinaga also related framed magnitude homology to open intervals determined by the frame. Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1. We also write $C_{*}(a,b)_{G}$ for the reduced chain complex of the order complex of $(a,b)_{G}$.
Proposition 2.23 ([13, Proposition 2.3]).
For $a,b\in V(G)$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
$MC_{*,\ell}(G;a,b)\to C_{*-2}(a,b)_{G},$
which induces an isomorphism of homology groups
$MH_{k,\ell}(G;a,b)\xrightarrow{\cong}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
Proof.
Any proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$ must be geodesic since $\ell=d(a,b)$. Mapping $\vec{x}$ to the linearly ordered chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism. ∎
Corollary 2.24.
The geodesic magnitude homology admits a decomposition
$MH_{k,\ell}^{\mathrm{geod}}(G)\cong\bigoplus_{\begin{subarray}{c}a,b\in V(G)\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$. A more general version of Proposition 2.23 is as follows.
Theorem 2.25 ([20, Theorem 4.4 and Corollary 4.5]).
Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$. Then there is an isomorphism of chain complexes
$MC^{\mathrm{gs}}_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G},$
which induces isomorphisms of homology groups
$MH^{\mathrm{gs}}_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
Remark 2.26.
The original statement of Theorem 4.4 in [20] does not contain the assumption $\ell(\vec{a})<m_{G}$. The result fails if a frame node $a_{i}$ of $\vec{a}$ can be smoothed out by nodes from neighboring intervals. For example, let $G=C_{4}$ be the cycle graph of $4$ vertices, labeled by $0,1,2,3$ in cyclic order. Choose $\vec{a}=(0,1,3)\in P_{2,3}(G)$. Then $(0,1)_{G}$ is empty and $(1,3)_{G}$ consists of two incomparable points $\{0,2\}$. Compare $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$ with $C_{-1}(\vec{a})_{G}$. The element $\varnothing\otimes 0$ in $C_{-1}(\vec{a})_{G}$ does have a preimage $(0,1,0,3)$ in $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$. However, the element $\varnothing\otimes 2$ in $C_{-1}(\vec{a})_{G}$ does not have a preimage because $(0,1,2,3)$ is a $4$-cut.
Proof of Theorem 2.25.
Let $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$ such that $\varphi(\vec{a})=\vec{a}$. Then $\ell=\ell(\vec{a})<m_{G}$. If $\vec{x}\in P_{k,\vec{a}}(G)$ is a basis element of $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$, then $\vec{x}$ has the form
$\vec{x}=(a_{0},x^{1}_{1},\ldots,x^{1}_{k_{1}},a_{1},x^{2}_{1},\ldots,x^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
Note that $k=m+\sum_{i=1}^{m}k_{i}$. The smooth nodes between $a_{i-1}$ and $a_{i}$ form a linearly ordered chain $\langle x^{i}\rangle:=(x^{i}_{1}\prec_{G}\cdots\prec_{G}x^{i}_{k_{i}})$ in $C_{k_{i}-1}(a_{i-1},a_{i})_{G}$. Define $\Phi_{k}:MC^{\mathrm{gs}}_{k,\vec{a}}(G)\to C_{k-2m}(\vec{a})_{G}$ by
$\Phi_{k}(\vec{x})=\langle x^{1}\rangle\otimes\langle x^{2}\rangle\otimes\cdots\otimes\langle x^{m}\rangle.$
The map $\Phi_{k}$ is injective on basis elements. To prove that $\Phi_{k}$ is also surjective, consider a basis element $\langle y^{1}\rangle\otimes\cdots\otimes\langle y^{m}\rangle$ of $C_{k-2m}(\vec{a})_{G}$, where $\langle y^{i}\rangle=(y^{i}_{1}\prec_{G}\cdots\prec_{G}y^{i}_{k_{i}})\in C_{k_{i}-1}(a_{i-1},a_{i})_{G}$ for $i=1,\ldots,m$. Let
$\vec{y}=(a_{0},y^{1}_{1},\ldots,y^{1}_{k_{1}},a_{1},y^{2}_{1},\ldots,y^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
We have $\vec{y}\in P_{k,\ell}(G)$. It remains to prove $\varphi(\vec{y})=\vec{a}$. First note that $(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i})$ is geodesic (2.1), and hence $y^{i}_{j}$ is smooth in $\vec{y}$. Next suppose, for contradiction, that $a_{i}$ is smooth in $\vec{y}$ for some $i=1,\ldots,m-1$, that is,
$d(y^{i}_{k_{i}},y^{i+1}_{1})=d(y^{i}_{k_{i}},a_{i})+d(a_{i},y^{i+1}_{1}).$
Here $y^{i}_{k_{i}}$ is understood as $a_{i-1}$ if $k_{i}=0$. Similarly $y^{i+1}_{1}$ is understood as $a_{i+1}$ if $k_{i+1}=0$. Note that $k_{i}$ and $k_{i+1}$ cannot both be zero, since otherwise $\varphi(\vec{a})\neq\vec{a}$. Consider the subchain
$\vec{z}=(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i},y^{i+1}_{1},\ldots,y^{i+1}_{k_{i+1}},a_{i+1}).$
It satisfies $\varphi(\vec{z})=(a_{i-1},a_{i+1})$ and $\ell(\vec{z})=d(a_{i-1},a_{i})+d(a_{i},a_{i+1})>d(a_{i-1},a_{i+1})$ by the assumption that $a_{i}$ is singular in $\vec{a}$. By Lemma 2.21, $\vec{z}$ contains a subchain $\vec{w}$ that is a $4$-cut. However, $\ell(\vec{w})\leq\ell(\vec{z})\leq\ell(\vec{y})=\ell$, which contradicts the assumption $\ell<m_{G}$. The verification that $\Phi_{k}$ defines a chain map is omitted. ∎
2.4. Hyperplane arrangements
The standard references are [28, 32]. Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{d}$. For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ for the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$. The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ in which two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane. The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$. Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, it is isomorphic to an isometric subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$.
We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusion. An element of $L(\mathcal{A})$ is called a flat. Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{d}$. The arrangement is called essential if the center is the origin. The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension: $\operatorname{rank}X=\operatorname{codim}X$. The rank of $\mathcal{A}$ is defined as $\operatorname{rank}\mathcal{A}=\operatorname{rank}(\cap\mathcal{A})$. We write $L_{k}(\mathcal{A})$ for the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ for the set of flats of rank at most $k$. Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$. It is classically known that $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},-1)|$ [35].
The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{d}$; the strata are called faces (covectors). Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusion. The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and the maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$. For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist. There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is, the minimal subspace of $\mathbb{R}^{d}$ containing $F$. We define the dimension of a face as the dimension of its support.
For a face $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$. Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$. Note that if $F,G\in\mathcal{F}(\mathcal{A})$ have the same support $X\in L(\mathcal{A})$, then $\mathcal{A}_{F}=\mathcal{A}_{G}=\mathcal{A}_{X}$.
Definition 2.27.
For a flat $X\in L(\mathcal{A})$, we denote by
$c^{X}:=\#s^{-1}(X)$
the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}=\mathbb{R}^{d}$, we also define the beta invariant of $X$ by
$\beta_{X}:=\beta(\mathcal{A}_{X}),$
where the beta invariant of an arrangement is defined as
$\beta(\mathcal{A})=(-1)^{\operatorname{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
Remark 2.28.
The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$. In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$. Then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (see [35, 2, 32]). Hence the product $c^{X}\beta_{X}$ is also a counting function:
$c^{X}\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid s(\overline{C}\cap H)=X\}.$
For an arrangement $\mathcal{A}$ in $V=\mathbb{R}^{d}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$. We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ for the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$. The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$. For a face $F\in\mathcal{F}(\mathcal{A})$, we say that the corresponding face $z_{F}$ of $Z(\mathcal{A})$ is dual to $F$. In particular, the 0-faces (vertices) of $Z(\mathcal{A})$ are dual to the chambers of $\mathcal{A}$.
With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\in\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$. Declaring $0<\pm$ determines a partial order on the set of sign vectors, which coincides with the face order of $\mathcal{F}(\mathcal{A})$. We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follows. For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(\mathcal{A})$ be the sign vector
$(FG)_{H}=\begin{cases}F_{H},\text{ if }F_{H}\neq 0;\\ G_{H},\text{ if }F_{H}=0.\end{cases}$
This product does not depend on the choice of defining forms: geometrically, $FG$ is the face reached by starting from a relative interior point of $F$ and moving a sufficiently small distance toward a relative interior point of $G$ (see Figure 1).
[…]
Figure 1. A geometric visualization of the non-commutative Tits product in the $A_{2}$ arrangement. Let $F$ be a ray (blue) and $G$ be a chamber (red). (a) The product FGFG represents taking a small step from $F$ toward $G$, which immediately lands in the adjacent chamber. (b) The product GFGF represents taking a small step from $G$ toward $F$. Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
2.5. Tope graphs and graph-theoretic properties
For an arrangement $\mathcal{A}$ in $\mathbb{R}^{d}$, note that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$. For a face $F\in\mathcal{F}(\mathcal{A})$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$; equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$. We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$. We first prove a useful lemma.
Lemma 2.29.
For $F\in\mathcal{F}(\mathcal{A})$, the subgraph $\mathcal{T}(\mathcal{A})_{F}$ of $\mathcal{T}(\mathcal{A})$ is gated with gate projection $D\mapsto FD$.
Proof.
We must show that the equality
(2.4) $d(D,C)=d(D,FD)+d(FD,C)$
holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$. In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (the number of unequal signs), where
$d(X,Y)_{H}=\begin{cases}1,\text{ if }X_{H}\neq Y_{H};\\ 0,\text{ if }X_{H}=Y_{H}.\end{cases}$
We prove that both sides of equation (2.4) have the same contribution on each $H\in\mathcal{A}$. If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$; also, since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$. Therefore (2.4) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$. Next suppose $H\in\mathcal{A}_{F}$. Then $F_{H}=0$ and $d(D,FD)_{H}=0$. The equation (2.4) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is true since $D_{H}=(FD)_{H}$. ∎
We record some important examples of tope graphs.
Example 2.30.
Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes in $\mathbb{R}^{d}$. Then the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
Example 2.31.
Let $r\geq 2$. An arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$ is said to be generic if any subarrangement of $r$ hyperplanes has rank $r$. Boolean arrangements are generic. Generic arrangements are realizations of uniform matroids $U_{r,n}$. By abuse of notation, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$. For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$. The tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ obtained by removing the all-++ vector and the all-- vector.
Example 2.32.
Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$. The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutahedron, and the tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the Cayley graph of $(W,S)$.
Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if it is a simplicial cone. An arrangement is simplicial if all its chambers are simplicial.
Proposition 2.33 ([4, Propositions 4.2.15 and 4.4.8]).
Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
  1. (1)
    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\operatorname{diam}(\mathcal{T})$.
  2. (2)
    The mapping $C\mapsto-C$ is a fixed-point-free automorphism of $\mathcal{T}$.
  3. (3)
    $\mathcal{T}$ is bipartite.
  4. (4)
    For each vertex $C$, we have $\deg(C)\geq r$, where equality holds if and only if $C$ is simplicial.
  5. (5)
    $\mathcal{T}$ is $r$-connected.
2.6. Varchenko matrix
Varchenko introduced a bilinear form associated with an arrangement and computed its determinant [33]. Generalizations to oriented matroids can be found in [18]. See [8] and [2, Sections 8.4–8.5] for details.
Here we only consider the equal-weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix by
$\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]_{C,D\in\mathrm{Ch}(\mathcal{A})}.$
In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). This matrix can also be defined for oriented matroids.
Theorem 2.34 ([33], see also [2, Theorem 8.23]).
The determinant of the $q$-Varchenko matrix is
$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}.$
2.7. Tope poset and generating function
Fix a chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$. This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$. The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by the rank function $\rho_{B}(C)=d(B,C)$. In the notation of Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
Proposition 2.35 ([9, Theorem 2.2], see also [4, Theorem 4.4.2]).
An open interval $(C_{1},C_{2})$ in $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\operatorname{rank}\mathcal{A}-\dim F-2}$ if $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise.
Corollary 2.36 ([10, Theorem 1.11], see also [4, Corollary 4.4.3]).
The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is given by
$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\operatorname{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}\text{ for some }F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
Proposition 2.37 ([3, Theorem 3.4]).
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A},B)$ is a lattice for any $B\in\mathrm{Ch}(\mathcal{A})$.
Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(B,C)}$ be the rank-generating function of $\mathcal{T}(\mathcal{A},B)$. Note that it is the sum of the entries in the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
Proposition 2.38 ([30]).
If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then we have
$D_{\mathcal{A},B}(q)=\prod_{i=1}^{\ell}[1+e_{i}]_{q}$
for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-number.
The generating-function decomposition formula has been generalized to supersolvable arrangements [3] and inductively factored arrangements [19]. See also [1, 27, 12].
3. Magnitude of arrangements
For a real central hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of its tope graph $\mathcal{T}(\mathcal{A})$.
3.1. Basic properties
We first prove the symmetry of the magnitude with respect to $q\leftrightarrow q^{-1}$.
Proposition 3.1.
Let $\mathcal{A}$ be an arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
$\mathrm{Mag}(\mathcal{A},q^{-1})=q^{N}\mathrm{Mag}(\mathcal{A},q).$
Proof.
Recall that there is an involution $\tau:\mathrm{Ch}(\mathcal{A})\to\mathrm{Ch}(\mathcal{A})$ sending $D$ to D-D. For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N. Thus the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
$q^{N}\mathcal{V}_{q^{-1}}=[q^{N-d(C,D)}]=[q^{d(C,-D)}]=\mathcal{V}_{q}P_{\tau},$
where $P_{\tau}$ is the permutation matrix induced by $\tau$. By the definition of magnitude, we have
$\mathrm{Mag}(\mathcal{A},q^{-1})=\mathbf{1}^{T}(\mathcal{V}_{q^{-1}})^{-1}\mathbf{1}=\mathbf{1}^{T}(q^{-N}\mathcal{V}_{q}P_{\tau})^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}P_{\tau}^{-1}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathrm{Mag}(\mathcal{A},q),$
where we used $\mathbf{1}^{T}P_{\tau}^{-1}=\mathbf{1}^{T}$. ∎
Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ satisfies the one-point property established in [22].
Proposition 3.2.
If $\mathcal{A}$ is an arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
Next we prove some structural results for $\mathrm{Mag}(\mathcal{A})$.
Theorem 3.3.
For an arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced expression of the magnitude, with the leading coefficient of $Q(q)$ positive. Then
  1. (1)
    $\mathcal{A}=\deg Q-\deg P$.
  2. (2)
    $P,Q$ are palindromic.
  3. (3)
    All roots of $Q$ are roots of unity other than $1$.
Proof.
Let $P(q)=a_{m}q^{m}+\cdots+a_{0}$ and $Q(q)=b_{n}q^{n}+\cdots+b_{0}$ where $a_{m}\neq 0$ and $b_{n}\neq 0$. Then $m=\deg P$ and $n=\deg Q$. Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$. By the symmetry formula (Proposition 3.1), we have
$q^{N}=\frac{\mathrm{Mag}(\mathcal{A},q^{-1})}{\mathrm{Mag}(\mathcal{A},q)}=\frac{P(q^{-1})Q(q)}{P(q)Q(q^{-1})}.$
Multiplying both sides by $q^{m-n}$, we obtain
$q^{N+m-n}=\frac{q^{m}P(q^{-1})Q(q)}{P(q)q^{n}Q(q^{-1})}=\frac{(a_{m}+\cdots+a_{0}q^{m})(b_{n}q^{n}+\cdots+b_{0})}{(a_{m}q^{m}+\cdots+a_{0})(b_{n}+\cdots+b_{0}q^{n})}.$
Comparing degrees, we obtain N+mn=0N+m-n=0, and (1) is proved.
Write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ for the reverse polynomials. We have proved
$\mathrm{Mag}(\mathcal{A},q)=\frac{P(q)}{Q(q)}=\frac{P^{*}(q)}{Q^{*}(q)}.$
Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced, since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, then $P$ and $Q$ would share the common root $1/\alpha$, contradicting the assumption that $P/Q$ is reduced. Hence we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$. Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
$Q^{*}(1)=Q(1)=rQ(1).$
Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, we have $Q(1)\neq 0$, so $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, and (2) is proved.
Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, the roots of $Q(q)$ are roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.34. We have already seen that $Q(1)\neq 0$. Thus (3) is proved. ∎
If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as the Cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$. Therefore, by Proposition 2.5, we have a decomposition of magnitude.
Proposition 3.4.
$\mathrm{Mag}(\mathcal{A}_{1}\oplus\mathcal{A}_{2})=\mathrm{Mag}(\mathcal{A}_{1})\mathrm{Mag}(\mathcal{A}_{2})$.
Coxeter arrangements carry a natural action of the corresponding Coxeter groups on the chambers, which makes their magnitudes easy to compute.
Proposition 3.5.
For a Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
$\mathrm{Mag}(\mathcal{A}_{W})=\frac{\#W}{\prod_{i}[1+e_{i}]_{q}}.$
Proof.
The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutahedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.32), and hence is vertex-transitive. The result follows from Proposition 2.3 and Proposition 2.38. ∎
3.2. Face decomposition formula
For reciprocity purposes, let us define the following.
Definition 3.6.
We define the interior magnitude of $\mathcal{A}$ by
$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\operatorname{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
where the latter equality follows from Proposition 3.1.
We prove the following “face decomposition formula” for the magnitude of an arrangement.
Theorem 3.7.
For any arrangement $\mathcal{A}$, we have
$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F}).$
Remark 3.8.
Theorem 3.7 can be reformulated as
$\mathrm{Mag}(\mathcal{A})=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\operatorname{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. It can also be written as
$\mathrm{Mag}(\mathcal{A})=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
This can be regarded as a recursive formula for computing $\mathrm{Mag}(\mathcal{A})$.
Remark 3.9.
Taking the limit $q\to 1$ gives $\mathrm{Mag}(\mathcal{A},1)=1$ and $\mathrm{Mag}^{\circ}(\mathcal{A},1)=(-1)^{\operatorname{rank}\mathcal{A}}$ by Proposition 3.2. Therefore, Theorem 3.7 can be regarded as a $q$-analogue of the decomposition of the Euler characteristic of the zonotope into its open faces.
Corollary 3.10.
The magnitude $\mathrm{Mag}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is determined by $L(\mathcal{A})$.
Lemma 3.11.
For any arrangement $\mathcal{A}$ and $D\in\mathrm{Ch}(\mathcal{A})$, we have
$\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}=1.$
Proof.
If $\operatorname{rank}\mathcal{A}=0$, then $\mathcal{A}$ is the empty arrangement and there is nothing to prove, so let us assume $r:=\operatorname{rank}\mathcal{A}\geq 1$. Passing to the essentialization does not change the chamber set, the tope graph, or the quantities $\mathcal{A}_{F}$ and $\operatorname{codim}F=\operatorname{rank}(\mathcal{A}_{F})$, so we may work in an essential arrangement of rank $r$.
Choose defining forms $\alpha_{H}$ so that $D_{H}=+$ for every $H\in\mathcal{A}$. For a face $F$, define $M_{D}(F)\subset\mathcal{A}$ by
$M_{D}(F):=\{H\in\mathcal{A}:F_{H}\neq+\}.$
Since $F_{H}=0$ exactly when $H\in\mathcal{A}_{F}$, while $d(D,FD)$ counts the hyperplanes on which $F_{H}=$-, we have $M_{D}(F)=\#\mathcal{A}_{F}+d(D,FD)$. Therefore it is enough to prove
$S_{D}(q):=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#M_{D}(F)}=1.$
Expand each power as
$q^{\#M_{D}(F)}=(1+(q-1))^{\#M_{D}(F)}=\sum_{T\subseteq M_{D}(F)}(q-1)^{\#T}.$
After exchanging the order of summation we obtain
$S_{D}(q)=\sum_{T\subseteq\mathcal{A}}(q-1)^{\#T}E_{T},\qquad E_{T}:=\sum_{F:\,T\subseteq M_{D}(F)}(-1)^{\operatorname{codim}F}.$
Hence the desired identity will follow once we show $E_{T}=\mathbf{1}_{T=\varnothing}$, i.e. $E_{T}=0$ if $T\neq\varnothing$ and $E_{\varnothing}=1$. We interpret $E_{T}$ as a reduced Euler characteristic. Let $S^{r-1}$ be the unit sphere in the ambient space $V=\mathbb{R}^{r}$. For each hyperplane $H\in\mathcal{A}$, let
$H_{D}^{-}:=\{x\in S^{r-1}:\alpha_{H}(x)\leq 0\}$
be the closed hemisphere opposite to $D$, and for $T\subseteq\mathcal{A}$ define
$\Sigma_{T}:=\bigcap_{H\in T}H_{D}^{-}\subseteq S^{r-1},\qquad\Sigma_{\varnothing}:=S^{r-1}.$
The arrangement induces a regular CW decomposition of $S^{r-1}$ where the open cells are $F\cap S^{r-1}$ for non-minimal faces $F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}$. An open cell $F\cap S^{r-1}$ lies in $\Sigma_{T}$ if and only if $T\subseteq M_{D}(F)$. Therefore
$\chi(\Sigma_{T})=\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{\dim F-1}.$
Since $\operatorname{codim}F=r-\dim F$ and the minimal face $\cap\mathcal{A}$ contributes $(-1)^{r}$ to $E_{T}$, we obtain
$E_{T}$ $=(-1)^{r}+\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{r-\dim F}=(-1)^{r-1}\left(\chi(\Sigma_{T})-1\right).$
If $T=\varnothing$, then $\Sigma_{\varnothing}=S^{r-1}$, so $E_{\varnothing}=(-1)^{r-1}(\chi(S^{r-1})-1)=1$. If $T\neq\varnothing$, then $\Sigma_{T}$ is the intersection of a nonempty collection of closed hemispheres, so $\Sigma_{T}$ is homeomorphic to $D^{r-1}$. Therefore $E_{T}=(-1)^{r-1}(\chi(D^{r-1})-1)=0$. This shows $E_{T}=\mathbf{1}_{T=\varnothing}$. ∎
Proof of Theorem 3.7.
Fix a face $F\in\mathcal{F}(\mathcal{A})$. Let $w_{F}\colon\mathrm{Ch}(\mathcal{A})_{F}\to\mathbb{Q}(q)$ be the weight vector of $\mathcal{T}(\mathcal{A})_{F}\cong\mathcal{T}(\mathcal{A}_{F})$, so that
$\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(E,C)}w_{F}(C)=1$
for every $E\in\mathrm{Ch}(\mathcal{A})_{F}$. Extend $w_{F}$ by zero outside $\mathrm{Ch}(\mathcal{A})_{F}$. Now define a function $\widetilde{w}_{\mathcal{A}}\colon\mathrm{Ch}(\mathcal{A})\to\mathbb{Q}(q)$ by
$\widetilde{w}_{\mathcal{A}}(C):=\sum_{F\leq C}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}w_{F}(C).$
Taking the sum over all chambers gives
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}w_{F}(C)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F},q)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
Therefore Theorem 3.7 will follow once we prove that $\widetilde{w}_{\mathcal{A}}$ is the actual weight vector of $\mathcal{A}$. Fix a chamber $D\in\mathrm{Ch}(\mathcal{A})$. We compute the $D$-th coordinate of $\mathcal{V}_{q}(\mathcal{A})\widetilde{w}_{\mathcal{A}}$:
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(D,C)}w_{F}(C).$
Since $\mathcal{T}(\mathcal{A})_{F}$ is gated with gate projection $D\mapsto FD$ (Lemma 2.29), we have
$d(D,C)=d(D,FD)+d(FD,C)$
for every $C\in\mathrm{Ch}(\mathcal{A})_{F}$. Using this identity and the defining equation for the local weight vector $w_{F}$, we obtain
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(FD,C)}w_{F}(C)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}.$
By Lemma 3.11, this equals $1$ for every $D\in\mathrm{Ch}(\mathcal{A})$. ∎
3.3. Examples
We use the notation $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [15].
Example 3.12.
The near-pencil $\mathcal{A}(n,0)$ has the tope graph $C_{2(n-1)}\Box K_{2}$ and hence
$\mathrm{Mag}(\mathcal{A}(n,0))=\frac{4(n-1)}{[2]_{q}^{2}[n-1]_{q}}.$
Example 3.13.
Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$. Therefore, the magnitudes of both arrangements can be computed from Proposition 3.5. In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
$\mathrm{Mag}(\mathcal{A}(6,1))$ $=\frac{24}{[4]_{q}!}=\frac{24}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)}$
$=24-72q+96q^{2}-72q^{3}+48q^{4}-72q^{5}+120q^{6}-144q^{7}+144q^{8}-144q^{9}+144q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(9,1))$ $=\frac{48}{[2]_{q}[4]_{q}[6]_{q}}=\frac{48}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)}$
$=48-144q+192q^{2}-192q^{3}+240q^{4}-336q^{5}+432q^{6}-528q^{7}+624q^{8}-720q^{9}+816q^{10}+\cdots$
Additional examples are given in the appendix.
4. Computations for rank 3 arrangements
4.1. Magnitude of rank 3 arrangements
Using the face decomposition formula (Theorem 3.7), we can explicitly compute the magnitude of rank $3$ arrangements.
Let $\mathcal{A}$ be an arrangement of rank $3$ with $n=\#\mathcal{A}$. For $k\geq 2$, we write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$.
Theorem 4.1.
Let $\mathcal{A}$ be an arrangement of rank $3$ with $\mathcal{A}=n$. Then we have
$\mathrm{Mag}(\mathcal{A})=\dfrac{1}{1+q^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}\biggr).$
Proof.
Let $\mathcal{F}_{d}(\mathcal{A})$ denote the set of faces of codimension $d$ of $\mathcal{A}$. By Theorem 3.7, we have
$\mathrm{Mag}(\mathcal{A})$ ${}=\sum_{F\in\mathcal{F}_{0}(\mathcal{A})}1-\sum_{F\in\mathcal{F}_{1}(\mathcal{A})}\dfrac{2q}{1+q}+\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}\dfrac{2\#\mathcal{A}_{F}q^{\#\mathcal{A}_{F}}(1-q)}{(1+q)(1-q^{\#\mathcal{A}_{F}})}-q^{n}\mathrm{Mag}(\mathcal{A},q)$
${}=\#\mathrm{Ch}(\mathcal{A})-\#\mathcal{F}_{1}(\mathcal{A})\cdot\dfrac{2q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}-q^{n}\mathrm{Mag}(\mathcal{A},q).$
For each $F\in\mathcal{F}_{2}(\mathcal{A})$, the number of faces $G\in\mathcal{F}_{1}(\mathcal{A})$ of codimension $1$ adjacent to $F$ is exactly $2\#\mathcal{A}_{F}$. Conversely, for each $G\in\mathcal{F}_{1}(\mathcal{A})$, there are exactly two faces $F\in\mathcal{F}_{2}(\mathcal{A})$ of codimension $2$ adjacent to $G$. Therefore we have
$\mathcal{F}_{1}(\mathcal{A})=\dfrac{1}{2}\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}2\#\mathcal{A}_{F}=\sum_{k\geq 2}kn_{k}.$
Using this identity, we obtain
$(1+q^{n})\mathrm{Mag}(\mathcal{A})$ ${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}$
${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}.$
This completes the proof. ∎
Example 4.2.
The arrangement $\mathcal{A}(6,1)$ from Grünbaum’s catalogue [15] satisfies
$\mathrm{Ch}(\mathcal{A}(6,1))=24\quad\text{and}\quad n_{k}=\begin{cases}6,&\text{ if }k=2;\\ 8,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
By Theorem 4.1, its magnitude is given by
$\mathrm{Mag}(\mathcal{A}(6,1))$ ${}=\dfrac{1}{1+q^{6}}\biggl(24-\dfrac{24q}{(1+q)^{2}}-\dfrac{48q}{1+q+q^{2}}\biggr)$
${}=\frac{24}{(1+q)^{2}(1+q+q^{2})(1+q^{2})}.$
This coincides with the formula given in Example 3.13.
4.2. Alternating sign property
In all the examples shown in Appendix A, the coefficients in the series expansion of the magnitude of each arrangement seem to have alternating signs. That is, if we write
$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell},$
then $(-1)^{\ell}c_{\ell}\geq 0$ appears to hold. Here we prove that if $\operatorname{rank}\mathcal{A}\leq 3$, then $(-1)^{\ell}c_{\ell}>0$ holds for all sufficiently large $\ell$. On the other hand, in the next subsection, we construct an example for which the inequality $(-1)^{\ell}c_{\ell}\geq 0$ fails for small $\ell$.
Definition 4.3.
For each $k\geq 2$, we define a rational function $G_{k}(q)$ and a sequence $(g_{k,\ell})_{\ell\geq 0}$ by
$G_{k}(q)=\dfrac{q(1-(-q)^{k-1})}{(1-q)(1-(-q)^{k})}=\sum_{\ell=0}^{\infty}g_{k,\ell}q^{\ell}.$
Theorem 4.1 implies that, for an arrangement $\mathcal{A}$ of rank $3$ with $\mathcal{A}=n$, we have
(4.1) $\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr).$
Lemma 4.4.
Let $\mathcal{A}$ be an arrangement of rank $3$. Let $n_{k}$ be the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$. Then we have $n_{2}>0$.
Proof.
We may assume that $\mathcal{A}$ is essential. Consider the zonotope $Z(\mathcal{A})$ of $\mathcal{A}$, which is a polytope in $\mathbb{R}^{3}$. The facets of $Z(\mathcal{A})$ are polygons with an even number of vertices. The number of facets with $2k$ vertices is exactly $n_{k}$. If $n_{2}=0$, then every facet of $Z(\mathcal{A})$ must have at least $6$ vertices. However, Euler’s polyhedral formula shows that this is impossible. Therefore we have $n_{2}>0$. ∎
Theorem 4.5.
Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}\leq 3$. Let $c_{\ell}$ denote the coefficient of $q^{\ell}$ in the series expansion of $\mathrm{Mag}(\mathcal{A})$:
$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}.$
Then we have $(-1)^{\ell}c_{\ell}>0$ for all sufficiently large $\ell$.
Proof.
When $\operatorname{rank}\mathcal{A}\leq 2$, the claim can be checked directly. Therefore, we may assume that $\operatorname{rank}\mathcal{A}=3$. Let $\mathcal{A}=n$. Write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$. Define $(e_{\ell})_{\ell\geq 0}$ by
$\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr)=\sum_{\ell=0}^{\infty}e_{\ell}q^{\ell}.$
It suffices to show that $e_{\ell}>0$ for all sufficiently large $\ell$. Put
$H(q):=\#\mathrm{Ch}(\mathcal{A})+\sum_{k=2}^{\infty}2kn_{k}G_{k}(q),\qquad H(q)=\sum_{\ell\geq 0}a_{\ell}q^{\ell}.$
We first record two elementary facts about the coefficients $g_{k,\ell}$ of $G_{k}$. If $k$ is even, then
$G_{k}(q)=\frac{q(1+q^{k-1})}{(1-q)(1-q^{k})}=\frac{q}{1-q}\sum_{j\geq 0}\left(q^{kj}+q^{kj+k-1}\right).$
Hence $g_{k,\ell}$ is non-negative and non-decreasing in $\ell$. In particular, for $k=2$ we have $g_{2,\ell}=\ell$. If $k$ is odd, then
$G_{k}(q)=\frac{q(1-q^{k-1})}{(1-q)(1+q^{k})}=\frac{q(1+q+\cdots+q^{k-2})}{1+q^{k}}.$
Thus the coefficient sequence $(g_{k,\ell})_{\ell\geq 0}$ is periodic and bounded. Since $n_{2}>0$ by Lemma 4.4, there is a constant $C>0$ such that $a_{\ell}\geq 4n_{2}\ell-C$ for all $\ell\geq 0$.
We now distinguish the parity of $n$. First suppose that $n$ is odd. Then $1+(-q)^{n}=1-q^{n}$, so
$\mathrm{Mag}(\mathcal{A},-q)=\frac{H(q)}{1-q^{n}}=H(q)\sum_{j\geq 0}q^{nj}.$
Therefore, with $N=\lfloor\ell/n\rfloor$,
$e_{\ell}=\sum_{j=0}^{N}a_{\ell-nj}\geq\sum_{j=0}^{N}\bigl(4n_{2}(\ell-nj)-C\bigr).$
The right-hand side is a polynomial of degree $1$ in $\ell$ with leading coefficient $4n_{2}(N+1)>0$, so it tends to $\infty$ as $\ell\to\infty$. Hence $e_{\ell}>0$ for all sufficiently large $\ell$ in this case.
Next suppose that $n$ is even. Then $1+(-q)^{n}=1+q^{n}$ and
$\frac{1}{1+q^{n}}=\sum_{j\geq 0}(-1)^{j}q^{nj}.$
For even $k$, let $b_{k,\ell}$ be the coefficient of $q^{\ell}$ in the series expansion of $\frac{G_{k}(q)}{1+q^{n}}$. Then we have
$b_{k,\ell}=\sum_{j=0}^{\lfloor\ell/n\rfloor}(-1)^{j}g_{k,\ell-nj}.$
Since $(g_{k,\ell})_{\ell}$ is non-negative and non-decreasing, pairing adjacent terms gives $b_{k,\ell}\geq 0$ for every $\ell$. For $k=2$, using $g_{2,m}=m$, we also get $b_{2,\ell}\to+\infty$. We claim that the odd $k$ terms, after division by $1+q^{n}$, are bounded. Indeed, for odd $k$ and even $n$,
$\frac{G_{k}(q)}{1+q^{n}}=\frac{q(1+q+\cdots+q^{k-2})}{(1+q^{k})(1+q^{n})}.$
Moreover $\gcd(1+q^{k},1+q^{n})=1$, because $k$ is odd and $n$ is even. Hence the denominator has only simple roots of unity, and the coefficient sequence is bounded. The coefficient sequence of $\mathrm{Ch}(\mathcal{A})/(1+q^{n})$ is also bounded. Thus there is a constant $D>0$ such that the total contribution of $\mathrm{Ch}(\mathcal{A})$ and all odd $k$ terms to $e_{\ell}$ is at least D-D for every $\ell$. Combining these observations, and using Lemma 4.4, we obtain
$e_{\ell}\geq 4n_{2}b_{2,\ell}-D.$
Since $b_{2,\ell}\to+\infty$, the right-hand side is positive for all sufficiently large $\ell$. This proves the theorem. ∎
4.3. Arrangements with non-alternating magnitude
In this section, we use the construction of Burr–Grünbaum–Sloane [5] for the orchard problem to construct an example of an arrangement whose magnitude does not have the alternating-sign property.
Recall the formula (4.1) for $\mathrm{Mag}(\mathcal{A},-q)$. The series expansion of $G_{3}(q)$ is given by
$G_{3}(q)=\dfrac{q}{1-q+q^{2}}=q+q^{2}-q^{4}-q^{5}+q^{7}+q^{8}-\cdots.$
This suggests that for an arrangement with large $n_{3}$, the series expansion of $\mathrm{Mag}(\mathcal{A},-q)$ might have negative coefficients. This motivates the construction given below.
Definition 4.6.
Let $S\subset\mathbb{R}^{2}$ be a finite set of points in $\mathbb{R}^{2}$. For each $k\geq 2$, define $m_{k}$ to be the number of lines that contain exactly $k$ points from $S$. Define an arrangement $\mathcal{A}(S)$ in $\mathbb{R}^{3}$ by
$\mathcal{A}(S)=\{(x,y,1)^{\perp}\mid(x,y)\in S\}.$
The arrangement $\mathcal{A}(S)$ is essential unless $S$ is contained in a single line. It satisfies $n_{k}=2m_{k}$ for $k\geq 2$.
We use the following result to construct an $S$ with large $m_{3}$.
Theorem 4.7 ([5, Theorem 1], see also [14, Proposition 2.6]).
For any $N\geq 3$, there exists a finite set $S\subset\mathbb{R}^{2}$ with #S=N\#S=N such that
$m_{k}=\begin{cases}N-1-2\cdot\mathbf{1}_{3\mid N},&\text{ if }k=2;\\ \biggl\lfloor\dfrac{N(N-3)}{6}+1\biggr\rfloor,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}$
where $\mathbf{1}_{3\mid N}$ is equal to $1$ when $3$ divides $N$ and 0 otherwise.
Example 4.8.
As an example of Theorem 4.7, we construct a set of points $S\subset\mathbb{R}^{2}$ with #S=18\#S=18 such that
$m_{k}=\begin{cases}15,&\text{ if }k=2;\\ 46,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
Let $X$ be the set of non-singular points of the real projective algebraic curve
$Y^{2}=X^{2}(X-Z).$
Then the points of $X$ naturally form a group isomorphic to $\mathbb{R}/\mathbb{Z}$. An explicit parametrization is given by
$\phi(t)=\bigl[\cot^{2}(\pi t)+1:\cot(\pi t)(\cot^{2}(\pi t)+1):1\bigr]\quad(t\in\mathbb{R}/\mathbb{Z}).$
Let $P_{j}\in\mathbb{R}^{2}$ be the image of $\phi(j/18)$ under the projective transformation $X:Y:Z]\mapsto[X:Y:X+Y+Z]$. Then $P_{i},P_{j},P_{k}$ lie on a single line if and only if $i+j+k\equiv 0\pmod{18}$. Define $S\subset\mathbb{R}^{2}$ by
$S=\{P_{j}\mid j=0,1,\dots,17\}.$
There are exactly $46$ triples $\{i,j,k\}\subset\mathbb{Z}/18\mathbb{Z}$ such that $i+j+k\equiv 0\pmod{18}$, so we have $m_{3}=46$. Also, $m_{k}=0$ for $k\geq 4$. Since $m_{2}+3m_{3}=\binom{18}{2}$ by double-counting, we have $m_{2}=15$.
Example 4.9.
We give an explicit construction of an arrangement $\mathcal{A}$ such that the series expansion of $\mathrm{Mag}(\mathcal{A})$ does not have alternating signs. Let $S\subset\mathbb{R}^{2}$ be a finite set as in Example 4.8. Then the associated arrangement $\mathcal{A}(S)$ satisfies
$\mathcal{A}(S)=18,\quad n_{k}=\begin{cases}30,&\text{ if }k=2;\\ 92,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}\quad\text{and}\quad\#\mathrm{Ch}(\mathcal{A})=216.$
Figure 2 shows the projectivization of $\mathcal{A}(S)$. By Theorem 4.1, the magnitude of $\mathcal{A}(S)$ is given by
$\mathrm{Mag}(\mathcal{A})$ ${}=\dfrac{1}{1+q^{18}}\biggl(216-\dfrac{120q}{(1+q)^{2}}-\dfrac{552q}{1+q+q^{2}}\biggr)$
${}=216-672q+792q^{2}-360q^{3}-72q^{4}-48q^{5}+720q^{6}-1392q^{7}+1512q^{8}-\cdots.$
In particular, the coefficient of $q^{4}$ is 72<0-72<0.
Figure 2. An arrangement whose magnitude does not have alternating signs
5. Magnitude homology of arrangements
We continue to denote invariants of $\mathcal{T}(\mathcal{A})$ simply by the corresponding invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}(\mathcal{A})$, and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$. Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
5.1. Small lengths
We begin our study of $MH_{k,\ell}(\mathcal{A})$ by investigating small lengths $\ell\leq 2$.
Proposition 5.1.
Let $\mathcal{A}$ be an arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
  1. (1)
    $MH_{0,0}(\mathcal{A})$ is free abelian of rank $\beta_{0,0}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A})$.
  2. (2)
    $MH_{1,1}(\mathcal{A})$ is free abelian of rank $\beta_{1,1}(\mathcal{A})=2\sum_{H\in\mathcal{A}}c^{H}$.
  3. (3)
    $MH_{2,2}(\mathcal{A})$ is free abelian of rank $\beta_{2,2}(\mathcal{A})=\beta_{1,1}(\mathcal{A})+4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$.
Proof.
(1) and (2) follow from (1) and (2) of Proposition 2.8, together with the observation that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of the restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.22 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$:
$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH^{\mathrm{gs}}_{2,\vec{a}}(\mathcal{A}).$
By Theorem 2.25, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets. For $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, where $a_{0},a_{1}$ are chambers at distance $2$, we may suppose $S(a_{0},a_{1})=\{H_{1},H_{2}\}$ and let $X=H_{1}\cap H_{2}\in L_{2}(\mathcal{A})$. If $\mathcal{A}_{X}=2$, then the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ consists of two incomparable chambers between $a_{0}$ and $a_{1}$. If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton. Since $MH^{\mathrm{gs}}_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology shows that only the first type of $(a_{0},a_{1})$, with $\mathcal{A}_{X}=2$, contributes a copy of $\mathbb{Z}$ to the magnitude homology. Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial. Thus we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent. The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty. Then $MH^{\mathrm{gs}}_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$. Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$. Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$. This proves (3).
To finish the proof, we need to show that all other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.22 again:
$MH_{1,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH^{\mathrm{gs}}_{1,\vec{a}}(\mathcal{A}).$
Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$. The interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty. Hence $MH^{\mathrm{gs}}_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$. ∎
5.2. Geodesic magnitude homology and diagonal tope graphs
We now determine the geodesic magnitude homology of arrangements. Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. Both $c^{X}$ and $c_{X}$ are determined by the intersection lattice $L(\mathcal{A})$.
Theorem 5.2.
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is given as follows:
$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ \operatorname{rank}\mathcal{A}_{F}=k,\ \#\mathcal{A}_{F}=\ell\end{subarray}}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}\cong\bigoplus_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}\mathbb{Z}^{c^{X}c_{X}}.$
In particular, $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
Proof.
By Corollary 2.24, we have
$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}a,b\in\mathrm{Ch}(\mathcal{A})\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
By Proposition 2.35, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\operatorname{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise. Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\operatorname{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contributes a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ to $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$, and there is no other summand. ∎
Corollary 5.3.
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude Betti numbers are given by
$\beta_{k,\ell}^{\mathrm{geod}}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}c^{X}c_{X}.$
As an application, we give a characterization of diagonal tope graphs.
Corollary 5.4.
Let $\mathcal{A}$ be an essential arrangement. Then $\mathcal{T}(\mathcal{A})$ is diagonal if and only if $\mathcal{A}$ is Boolean.
Proof.
We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14). Suppose $\mathcal{A}$ is not Boolean. Then $n=\#\mathcal{A}>r=\operatorname{rank}\mathcal{A}$. We have
$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{\mathrm{geod}}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
where the equality follows from Corollary 5.3. This means that $\mathcal{A}$ is not diagonal. ∎
5.3. Magnitude homology with support
Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, and define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$. Observe that $S(\vec{x})\leq\ell(\vec{x})$. One verifies that $x_{i}$ is smooth in $\vec{x}$, i.e. $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, if and only if $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$. Thus, if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$. In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes. This motivates the following definition.
Definition 5.5.
For an arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$. The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$. Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
The following result is a direct consequence of the above observation that the magnitude boundary operator preserves crossed hyperplanes.
Proposition 5.6.
For an arrangement $\mathcal{A}$, we have the following decomposition of magnitude chain groups:
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}),$
which induces a decomposition of homology groups
$MH_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MH_{k,\ell}(\mathcal{A};\mathcal{B}).$
Furthermore, $MC_{k,\ell}(\mathcal{A};\mathcal{B})=MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
Using the notion of magnitude homology with support, we define interior magnitude homology. Later, we show that this is a categorification of the interior magnitude defined in Definition 3.6.
Definition 5.7.
Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. We define the $X$-supported magnitude complex of $\mathcal{A}$ by
$MC_{*,\ell}(\mathcal{A};X):=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A};\mathcal{B}).$
Its homology $MH_{*,\ell}(\mathcal{A};X)$ is called the $X$-supported magnitude homology of $\mathcal{A}$. In particular, we define
$MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\cap\mathcal{A}),$
and call it the length $\ell$ interior magnitude complex of $\mathcal{A}$. We call $MH_{*,\ell}^{\circ}(\mathcal{A}):=H_{*}(MC_{*,\ell}^{\circ}(\mathcal{A}))$ the length $\ell$ interior magnitude homology of $\mathcal{A}$. We also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
5.4. Face decomposition for magnitude homology
Here we prove a homological version of the face decomposition formula established in Theorem 3.7. To this end, we introduce another notation.
Definition 5.8.
Let $\mathcal{A}$ be an arrangement and let $\mathcal{B}\subseteq\mathcal{A}$ be a subarrangement. Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$. We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$:
$P_{k,\ell}(G;\mathcal{B})=\{(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A};\mathcal{B})\mid x_{0},\ldots,x_{k}\in V(G)\}.$
Lemma 5.9.
For $F\in\mathcal{F}(\mathcal{A})$, the chain groups $MC_{*,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$ form a subcomplex of $MC_{*,\ell}(\mathcal{A};\mathcal{B})$.
Proof.
We must show that the boundary operator $\partial$ stabilizes the subgroup. By Proposition 5.6, $\partial$ strictly preserves the set $\mathcal{B}$ of crossed hyperplanes. Furthermore, by Lemma 2.29, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2). Because the boundary operator $\partial_{i}$ only deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained in the vertex set of $\mathcal{T}(\mathcal{A})_{F}$. Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$. ∎
Proposition 5.10.
Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. Then the $X$-supported magnitude complex can be decomposed as follows:
$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{*,\ell}^{\circ}(\mathcal{A}_{F}).$
Proof.
By Definition 5.7, $MC_{*,\ell}(\mathcal{A};X)=\bigoplus_{\mathcal{B}\subseteq\mathcal{A},\cap\mathcal{B}=X}MC_{*,\ell}(\mathcal{A};\mathcal{B})$. For a subarrangement $\mathcal{B}$ with $\cap\mathcal{B}=X$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$. By Definition 5.5, $S(\vec{x})=\mathcal{B}$. The chain $\vec{x}$ never crosses any hyperplane in $\mathcal{A}\setminus\mathcal{A}_{X}$. Therefore, every chamber in $\vec{x}$ has a constant nonzero sign on each hyperplane $H\in\mathcal{A}\setminus\mathcal{A}_{X}$. These signs determine a chamber $F\in\mathrm{Ch}(\mathcal{A}^{X})$, or equivalently a face $F\in\mathcal{F}(\mathcal{A})$ with $s(F)=X$. Therefore the chain $\vec{x}$ lies in the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$. By Lemma 2.29, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$ isometrically isomorphic to $\mathcal{T}(\mathcal{A}_{F})$. This induces a direct-sum decomposition of the $\mathcal{B}$-supported chains over the faces supported on $X$:
$MC_{k,\ell}(\mathcal{A};\mathcal{B})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}(\mathcal{A}_{F};\mathcal{B}).$
Summing both sides over $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=X$, we obtain
$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}_{F}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A}_{F};\mathcal{B}).$
Since we are summing over $F$ supported on $X$, we have $\mathcal{A}_{F}=\mathcal{A}_{X}$ and $\cap\mathcal{A}_{F}=X$, so the inner sum is exactly $MC_{*,\ell}^{\circ}(\mathcal{A}_{F})$. This proves the desired formula. ∎
We are now ready to state the main structural result for the magnitude homology of an arrangement.
Theorem 5.11.
For an arrangement $\mathcal{A}$, we have the following face decomposition of its magnitude complex:
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}),$
which induces the following face decomposition of magnitude homology:
$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
Proof.
By Proposition 5.6, the magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}).$
Grouping the $\mathcal{B}$-summands according to $\cap\mathcal{B}$, we have
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{X\in L(\mathcal{A})}MC_{k,\ell}(\mathcal{A};X).$
By Proposition 5.10, the $X$-supported complex decomposes into interior complexes
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}^{\circ}(\mathcal{A}_{F})=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
The decomposition of homology follows immediately. ∎
Corollary 5.12.
Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Then we have
$\mathrm{Mag}^{\circ}(\mathcal{A})=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
Moreover, we have $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$, and $\chi_{\ell}^{\circ}(\mathcal{A})=0$ for $\ell<n$.
Proof.
Define $\Delta(\mathcal{A})$ to be the difference of the two expressions:
$\Delta(\mathcal{A})=\mathrm{Mag}^{\circ}(\mathcal{A})-\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
We prove $\Delta(\mathcal{A})=0$ by induction on the rank of $\mathcal{A}$. By Theorem 3.7 and Theorem 5.11, we have
$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F}),\quad\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}_{\ell}(\mathcal{A}_{F}).$
Since $\mathrm{Mag}(\mathcal{A})$ can be written as $\sum_{\ell}\chi_{\ell}(\mathcal{A})q^{\ell}$, we obtain
$\sum_{F\in\mathcal{F}(\mathcal{A})}\Delta(\mathcal{A}_{F})=0.$
The induction hypothesis shows that $\Delta(\mathcal{A}_{F})=0$ for every face $F$ other than $\bigcap\mathcal{A}$. Thus $\Delta(\mathcal{A})=0$. The final assertion follows from the identity $\mathrm{Mag}^{\circ}(\mathcal{A})=(-1)^{r}q^{n}\mathrm{Mag}(\mathcal{A})$. ∎
Remark 5.13.
The identity $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ in Corollary 5.12 is analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31].
Remark 5.14.
In light of Corollary 5.12, it is natural to expect that there exists an isomorphism of homology groups
$MH_{k,\ell}(\mathcal{A})\cong MH_{k+r,\ell+n}^{\circ}(\mathcal{A})$
for all $k,\ell\geq 0$, and that the interior magnitude homology $MH_{k,\ell}^{\circ}(\mathcal{A})$ vanishes for $0\leq\ell<n$.
5.5. Magnitude homology on the diagonal
In this subsection, we give a combinatorial formula for the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$. First, we determine the interior magnitude homology of Boolean arrangements using Theorem 5.11.
Corollary 5.15.
For the Boolean arrangement $Bl_{d}$, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is concentrated on the diagonal $k=\ell$, and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
$\beta_{\ell,\ell}^{\circ}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$.
Proof.
We know from Corollary 5.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$. By Theorem 5.11, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence is concentrated on the diagonal $k=\ell$. Theorem 5.11 also gives
$\beta_{\ell,\ell}(Bl_{d})=\sum_{F\in\mathcal{F}(Bl_{d})}\beta_{\ell,\ell}^{\circ}((Bl_{d})_{F}).$
If $F\in\mathcal{F}(Bl_{d})$ has $\operatorname{codim}F=j$, then the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$. Applying Möbius inversion on $\mathcal{F}(Bl_{d})$ yields the desired formula. ∎
Next we prove the following vanishing result for diagonal interior magnitude homology of non-Boolean arrangements.
Theorem 5.16.
Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Suppose $\mathcal{A}$ is not Boolean, so that $r<n$. Then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes in all lengths:
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\quad\forall\ell\geq 0.$
Proof.
By Definition 5.7, we have
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=\cap\mathcal{A}\end{subarray}}MH_{\ell,\ell}(\mathcal{A};\mathcal{B}).$
We shall prove that each summand vanishes. In what follows, we suppose that $\mathcal{A}$ is essential without loss of generality, because passing to the essentialization does not alter the tope graph or the intersection lattice.
For $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=\cap\mathcal{A}=\{0\}$, let $m=\#\mathcal{B}$. Then $m\geq r$. Let $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ be the embedding of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{\mathcal{A}}$. Since $\phi$ is distance preserving, it induces an injective chain map $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{\mathcal{A}})$.
Let $c\in MC_{\ell,\ell}(\mathcal{A};\mathcal{B})$ be a cycle. It maps to a cycle $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{\mathcal{A}})$. By the endpoint decomposition (2.2), we may suppose without loss of generality that $c$ is a linear combination of $\mathcal{B}$-supported walks with a common starting vertex and a common terminal vertex. This means that the sequence of chambers in each walk of $c$ flips exactly the $m$ coordinates corresponding to $\mathcal{B}$, while the signs on the remaining nmn-m coordinates corresponding to $\mathcal{A}\setminus\mathcal{B}$ remain equal to those of the starting vertex. Geometrically, this means that the entire cycle $\phi_{\#}(c)$ is confined to an $m$-dimensional face of $Q_{\mathcal{A}}$, denoted by $Q_{\mathcal{B}}$ by abuse of notation.
We claim that this subcube $Q_{\mathcal{B}}$ contains a vertex $v_{0}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$. Indeed, in the case $m>r$, the hyperplanes of $\mathcal{B}$ contain linear dependencies. Consequently, the total number of geometrically realizable sign patterns on $\mathcal{B}$ is less than $2^{m}$. However, the subcube $Q_{\mathcal{B}}$ contains all $2^{m}$ sign patterns. Therefore, at least one sign pattern in this subcube is not geometrically realizable, so at least one vertex $v_{0}$ inside the subcube $Q_{\mathcal{B}}$ does not correspond to a chamber in $\mathcal{T}(\mathcal{A})$. In the case $m=r$, $\mathcal{B}$ is a Boolean subarrangement. Its $2^{r}$ chambers form full-dimensional orthants. Because $\mathcal{A}$ is not Boolean, there exists a hyperplane $H\in\mathcal{A}\setminus\mathcal{B}$, and we must have $r\geq 2$. This hyperplane $H$ cannot intersect the interior of every orthant. Thus, there exists a pair of antipodal chambers $C_{v},C_{-v}\in\mathrm{Ch}(\mathcal{B})$ separated by $H$ (that is, $H$ does not cut their interiors), where $v\in\{+,-\}^{\mathcal{B}}$ is the sign vector of $C_{v}$. We may assume that the extension of $v$ to $\tilde{v}\in\{+,-\}^{\mathcal{A}}$ using the signs of the starting vertex of $c$, regarded as a vertex of $Q_{\mathcal{A}}$, is geometrically realizable, i.e. is a vertex of $\mathcal{T}(\mathcal{A})$. Then the extension of v-v to $\widetilde{-v}\in\{+,-\}^{\mathcal{A}}$ is not geometrically realizable, since by definition $\widetilde{-v}_{H}=\widetilde{v}_{H}$, contradicting the geometric condition that $C_{v}$ and $C_{-v}$ are separated by $H$. Thus, we have shown the existence of such a forbidden vertex $v_{0}$ in the subcube $Q_{\mathcal{B}}$.
For a fixed $\ell$, $MC_{\ell+1,\ell}(Q_{\mathcal{A}})=0$; by abuse of notation, we regard $\phi_{\#}(c)$ as a homology class in $MH_{\ell,\ell}(Q_{\mathcal{A}})$. Recall that the isomorphism given in Example 2.14
$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{\mathcal{A}}),$
is induced by the Eilenberg–Zilber chain map [17]. We may assume that there are basis elements $b^{\lambda}$ of the left-hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$. We must show that $d_{\lambda}=0$.
Since $c$ crosses all hyperplanes in $\mathcal{B}$, its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{\mathcal{A}})$ is a linear combination of walks in $Q_{\mathcal{A}}$ along which every $\mathcal{B}$-coordinate is flipped at least once. This means that the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$. We write such an element as $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$. Thus $b^{\lambda}$ consists of $\ell$ steps in total in these $n$ dimensions with distribution $\lambda$.
By definition of the Eilenberg–Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps. It is a linear combination of walks in $Q_{\mathcal{A}}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminal vertex in $Q_{1}$, respectively. Since all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$, any vertex in the subcube $Q_{\mathcal{B}}$ must be visited by at least one walk in $\nabla(b^{\lambda})$. The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk because their step distributions differ.
$G^{+}(+,+,+)$$G^{-}(-,-,-)$(+,,)(+,-,-)Start: (+,+,)(+,+,-)(,+,)(-,+,-)(,+,+)(-,+,+)(,,+)(-,-,+)End: (+,,+)(+,-,+)Step 1 ($Z_{1}$)Step 2 ($X_{1}$)Step 3 ($X_{2}$)Step 4 ($Y_{1}$)
Figure 3. The tope graph of the $A_{2}$ arrangement embedded as a $C_{6}$ hexagon (blue/bold) inside the Boolean $Q_{3}$ cube. Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$. A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown passing through the forbidden vertex $G^{+}$. Note also that this walk cannot appear in any other $\nabla(b^{\mu})$.
Now we apply the above observations to the forbidden vertex $v_{0}$. Each $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 3), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$. We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$. Therefore $c=0$ and $MH_{\ell,\ell}(\mathcal{A};\mathcal{B})=0$. ∎
Remark 5.17.
By Corollary 5.15 and Theorem 5.16, we see that $\mathcal{A}$ is Boolean if and only if $\beta_{\ell,\ell}^{\circ}(\mathcal{A})\neq 0$ for some $\ell\geq 1$.
Corollary 5.18.
Let $\mathcal{A}$ be an arrangement. The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$. More precisely, for $\ell\geq 1$,
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$, where
$d_{\max}=\max\{\operatorname{rank}X\mid X\in L(\mathcal{A}),\ \operatorname{rank}X=\#\mathcal{A}_{X}\}.$
Proof.
Since $MC_{\ell+1,\ell}(\mathcal{A})=0$, the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian. By Theorem 5.11, the global magnitude homology admits a face decomposition. Restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F}).$
By Theorem 5.16, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e. $\mathcal{A}_{F}=\operatorname{rank}\mathcal{A}_{F}$. Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\operatorname{rank}X=\#\mathcal{A}_{X}$. For these faces, $\mathcal{A}_{F}\cong Bl_{\operatorname{rank}X}$. We can group the nonzero terms of the summation by their support flats $X\in L(\mathcal{A})$. For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$. Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 5.15), namely $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\operatorname{rank}X$, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
For $\ell\geq\operatorname{rank}X$, the binomial $\binom{\ell-1}{\operatorname{rank}X-1}$ is a polynomial in $\ell$ of degree $\operatorname{rank}X-1$. Therefore, for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$. ∎
Remark 5.19.
Example 2.14 and Proposition 5.1 (2)(3) are special cases of the formula in Corollary 5.18.
6. Open problems
Based on observations from examples in the appendix, we propose the following conjectures.
Conjecture 6.1.
Let $\mathcal{A}$ be an arrangement of rank $r$ and $n=\#\mathcal{A}$. Let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form. Then
  1. (1)
    Suppose $r\geq 3$ is odd. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{2n}(q)$ divides $Q(q)$.
  2. (2)
    Suppose $r\geq 4$ is even. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{n}(q)$ divides $Q(q)$.
  3. (3)
    $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0$ for sufficiently large $\ell$.
  4. (4)
    $MH_{k,\ell}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
  5. (5)
    $MH_{k,\ell}(\mathcal{A})$ is torsion-free.
  6. (6)
    Suppose that $r\geq 2$ and that $\mathcal{A}$ is indecomposable, that is, $\mathcal{A}$ cannot be expressed as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$ of nonzero arrangements. Then $\beta_{0,0}=\beta_{r,n}$.
Acknowledgements
The authors thank Weili Guo, Yoh Kitajima, Takuya Saito, and Masahiko Yoshinaga for their comments on a preliminary version of this paper.
Appendix A Computer experiments
The following examples are computed with SageMath [29].
Example A.1.
We list the magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [15] for $n\leq 13$. Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.12 and 3.13. We use the database of [6] to obtain the list of normal vectors.
$\mathrm{Mag}(\mathcal{A}(7,1))$ $=\frac{32q^{4}-40q^{2}+32}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{[2]_{q}^{2}[3]_{q}(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{14}(q)}$
$=32-96q+120q^{2}-72q^{3}+24q^{4}-48q^{5}+144q^{6}-272q^{7}+360q^{8}-336q^{9}+240q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(8,1))$ $=\frac{40q^{6}-8q^{4}-16q^{3}-8q^{2}+40}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{[4]_{q}!(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{16}(q)}$
$=40-120q+152q^{2}-112q^{3}+88q^{4}-136q^{5}+240q^{6}-344q^{7}+352q^{8}-248q^{9}+176q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,1))$ $=\frac{60q^{8}+60q^{7}-20q^{6}-40q^{5}-40q^{3}-20q^{2}+60q+60}{(q^{8}+4q^{7}+8q^{6}+11q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1)(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{[2]_{q}^{2}[3]_{q}[5]_{q}(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{20}(q)}$
$=60-180q+220q^{2}-140q^{3}+60q^{4}-80q^{5}+220q^{6}-380q^{7}+420q^{8}-340q^{9}+220q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,2))$ $=\mathrm{Mag}(\mathcal{A}(10,3))=\frac{60q^{6}-12q^{4}-48q^{3}-12q^{2}+60}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{[4]_{q}!(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)^{2}\Phi_{20}(q)}$
$=60-180q+228q^{2}-192q^{3}+204q^{4}-300q^{5}+432q^{6}-564q^{7}+660q^{8}-672q^{9}+576q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(11,1))$ $=\frac{72q^{6}-16q^{4}-64q^{3}-16q^{2}+72}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{[4]_{q}!(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{22}(q)}$
$=72-216q+272q^{2}-232q^{3}+256q^{4}-376q^{5}+528q^{6}-680q^{7}+800q^{8}-824q^{9}+784q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,1))$ $=\frac{84q^{6}-84q^{5}-36q^{4}+96q^{3}-36q^{2}-84q+84}{(q^{6}+2q^{5}+2q^{4}+2q^{3}+2q^{2}+2q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{[2]_{q}[6]_{q}(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{6}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+300q^{2}-168q^{3}+36q^{4}-84q^{5}+336q^{6}-588q^{7}+636q^{8}-504q^{9}+372q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,2))$ $=\frac{84q^{10}+84q^{9}+64q^{8}-4q^{7}-44q^{6}-128q^{5}-44q^{4}-4q^{3}+64q^{2}+84q+84}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{[5]_{q}!(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+316q^{2}-260q^{3}+252q^{4}-344q^{5}+508q^{6}-692q^{7}+804q^{8}-796q^{9}+760q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,3))$ $=\frac{84q^{6}-12q^{4}-96q^{3}-12q^{2}+84}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{[4]_{q}!(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+324q^{2}-312q^{3}+396q^{4}-564q^{5}+720q^{6}-876q^{7}+1044q^{8}-1128q^{9}+1116q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,1))$ $=\frac{96q^{8}-96q^{7}+72q^{6}-48q^{5}-48q^{3}+72q^{2}-96q+96}{(q^{8}+2q^{7}+3q^{6}+4q^{5}+4q^{4}+4q^{3}+3q^{2}+2q^{1}+1)(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{[2]_{q}^{2}(1+2q^{2}+2q^{4}+q^{6})(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)\Phi_{26}(q)}$
$=96-288q+360q^{2}-288q^{3}+264q^{4}-384q^{5}+624q^{6}-864q^{7}+984q^{8}-960q^{9}+888q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,2))$ $=\frac{96q^{6}-144q^{3}+96}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{[4]_{q}!(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=96-288q+384q^{2}-432q^{3}+624q^{4}-864q^{5}+1008q^{6}-1152q^{7}+1392q^{8}-1584q^{9}+1632q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,3))$ $=\frac{96q^{10}+96q^{9}+80q^{8}-8q^{7}-64q^{6}-160q^{5}-64q^{4}-8q^{3}+80q^{2}+96q+96}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{[5]_{q}!(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{26}(q)}$
$=96-288q+368q^{2}-328q^{3}+336q^{4}-424q^{5}+584q^{6}-784q^{7}+912q^{8}-920q^{9}+920q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,4))$ $=\frac{104q^{6}-56q^{4}-48q^{3}-56q^{2}+104}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{[4]_{q}!(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=104-312q+360q^{2}-192q^{3}+72q^{4}-168q^{5}+432q^{6}-696q^{7}+792q^{8}-672q^{9}+504q^{10}+\cdots$
Example A.2.
Consider the rank $3$ generic arrangement $\mathcal{A}=U_{3,4}$ defined by $xyz(x+y+z)$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{14q^{2}-20q+14}{(q^{2}+2q+1)(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{[2]_{q}^{2}(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{\Phi_{2}(q)^{2}\Phi_{8}(q)}$
$=14-48q+96q^{2}-144q^{3}+178q^{4}-192q^{5}+192q^{6}-192q^{7}+206q^{8}-240q^{9}+288q^{10}+\cdots$
Example A.3.
Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{18q^{4}-2q^{3}-8q^{2}-2q+18}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{[2]_{q}^{2}[3]_{q}(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{10}(q)}$
$=18-56q+88q^{2}-96q^{3}+104q^{4}-154q^{5}+248q^{6}-336q^{7}+376q^{8}-392q^{9}+450q^{10}+\cdots$
Example A.4.
Consider the graphic arrangement $\mathcal{A}$ of the graph $K_{5}\setminus e$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{[2]_{q}[4]_{q}[9]_{q}(q^{5}+1)}$
$=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{9}(q)\Phi_{10}(q)}$
$=96-396q+756q^{2}-924q^{3}+996q^{4}-1320q^{5}+1956q^{6}-2652q^{7}+3252q^{8}-3756q^{9}+4212q^{10}+\cdots$
where the numerator is $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator is $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
Example A.5.
Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]).
$\mathrm{Mag}(\mathcal{A})$ $=\frac{6(17q^{8}-4q^{7}-8q^{6}-9q^{5}+44q^{4}-9q^{3}-8q^{2}-4q+17)}{\Phi_{2}(q)^{3}\Phi_{3}(q)^{2}\Phi_{9}(q)\Phi_{10}(q)}$
$=102-432q+864q^{2}-1176q^{3}+1584q^{4}-2628q^{5}+4344q^{6}-6048q^{7}+7200q^{8}-8130q^{9}+9684q^{10}+\cdots$
where the numerator is $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator is $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
Example A.6.
We collect some results here on magnitude Betti numbers computed with the SageMath code of [17].
[…]
0 4 0 0 0 0 0 0 0 0
1 0 8 0 0 0 0 0 0 0
2 0 0 12 0 0 0 0 0 0
3 0 0 0 16 0 0 0 0 0
4 0 0 0 0 20 0 0 0 0
5 0 0 0 0 0 24 0 0 0
6 0 0 0 0 0 0 28 0 0
7 0 0 0 0 0 0 0 32 0
8 0 0 0 0 0 0 0 0 36
Table 1. Magnitude Betti numbers $\beta_{k,\ell}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
[…]
0 6 0 0 0 0 0 0 0 0
1 0 12 0 0 0 0 0 0 0
2 0 0 12 6 0 0 0 0 0
3 0 0 0 12 12 0 0 0 0
4 0 0 0 0 12 12 6 0 0
5 0 0 0 0 0 12 12 12 0
6 0 0 0 0 0 0 12 12 12
7 0 0 0 0 0 0 0 12 12
8 0 0 0 0 0 0 0 0 12
Table 2. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
[…]
0 8 0 0 0 0 0 0 0 0
1 0 16 0 0 0 0 0 0 0
2 0 0 16 0 8 0 0 0 0
3 0 0 0 16 0 16 0 0 0
4 0 0 0 0 16 0 16 0 8
5 0 0 0 0 0 16 0 16 0
6 0 0 0 0 0 0 16 0 16
7 0 0 0 0 0 0 0 16 0
8 0 0 0 0 0 0 0 0 16
Table 3. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
[…]
0 24 0 0 0 0 0 0 0 0
1 0 72 0 0 0 0 0 0 0
2 0 0 96 48 0 0 0 0 0
3 0 0 0 120 96 0 24 0 0
4 0 0 0 0 144 96 48 72 0
5 0 0 0 0 0 168 96 96 96
6 0 0 0 0 0 0 192 96 96
7 0 0 0 0 0 0 0 216 96
8 0 0 0 0 0 0 0 0 240
Table 4. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{3}$.
[…]
0 120 0 0 0 0 0 0
1 0 480 0 0 0 0 0
2 0 0 840 360 0 0 0
3 0 0 0 1200 960 0 240
4 0 0 0 0 1560 1440 360
5 0 0 0 0 0 1920 1920
6 0 0 0 0 0 0 2280
Table 5. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{4}$.
[…]
0 14 0 0 0 0 0 0 0 0
1 0 48 0 0 0 0 0 0 0
2 0 0 96 0 0 0 0 0 0
3 0 0 0 144 14 0 0 0 0
4 0 0 0 0 192 48 0 0 0
5 0 0 0 0 0 240 96 0 0
6 0 0 0 0 0 0 288 144 14
7 0 0 0 0 0 0 0 336 192
8 0 0 0 0 0 0 0 0 384
Table 6. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$; see Example A.2.
[…]
0 30 0 0 0 0 0 0
1 0 140 0 0 0 0 0
2 0 0 380 0 0 0 0
3 0 0 0 780 0 0 0
4 0 0 0 0 1340 30 0
5 0 0 0 0 0 2060 140
6 0 0 0 0 0 0 2940
Table 7. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{4,5}$.
[…]
0 18 0 0 0 0 0 0 0 0
1 0 56 0 0 0 0 0 0 0
2 0 0 88 24 0 0 0 0 0
3 0 0 0 120 48 18 0 0 0
4 0 0 0 0 152 48 80 0 0
5 0 0 0 0 0 184 48 136 24
6 0 0 0 0 0 0 216 48 168
7 0 0 0 0 0 0 0 248 48
8 0 0 0 0 0 0 0 0 280
Table 8. Magnitude Betti numbers $\beta_{k,\ell}$ for the graphic arrangement of $K_{4}\setminus e$; see Example A.3.
[…]
0 32 0 0 0 0 0 0 0 0
1 0 96 0 0 0 0 0 0 0
2 0 0 120 72 0 0 0 0 0
3 0 0 0 144 144 0 0 32 0
4 0 0 0 0 168 144 72 0 96
5 0 0 0 0 0 192 144 144 0
6 0 0 0 0 0 0 216 144 144
7 0 0 0 0 0 0 0 240 144
8 0 0 0 0 0 0 0 0 264
Table 9. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(7,1)$
[…]
0 40 0 0 0 0 0 0 0 0
1 0 120 0 0 0 0 0 0 0
2 0 0 152 72 16 0 0 0 0
3 0 0 0 184 144 32 0 0 40
4 0 0 0 0 216 144 104 0 16
5 0 0 0 0 0 248 144 176 0
6 0 0 0 0 0 0 280 144 176
7 0 0 0 0 0 0 0 312 144
8 0 0 0 0 0 0 0 0 344
Table 10. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(8,1)$
[…]
0 48 0 0 0 0 0 0 0 0
1 0 144 0 0 0 0 0 0 0
2 0 0 192 48 48 0 0 0 0
3 0 0 0 240 96 96 0 0 0
4 0 0 0 0 288 96 144 0 48
5 0 0 0 0 0 336 96 192 0
6 0 0 0 0 0 0 384 96 192
7 0 0 0 0 0 0 0 432 96
8 0 0 0 0 0 0 0 0 480
Table 11. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(9,1)$